### v2

# The spherical maximal operators on hyperbolic spaces

Abstract.
In this article we investigate $L^{p}$ boundedness of the spherical maximal operator $\mathfrak{m}^{\alpha}$ of (complex) order $\alpha$ on the $n$-dimensional hyperbolic space $\mathbb{H}^{n}$, which was introduced and studied by Kohen [13]. We prove that when $n\geq 2$, for $\alpha\in\mathbb{R}$ and $1<p<\infty$, if
$\|\mathfrak{m}^{\alpha}(f)\|_{L^{p}(\mathbb{H}^{n})}\leq C\|f\|_{L^{p}(\mathbb{H}^{n})},$
then we must have $\alpha>1-n+n/p$ for $1<p\leq 2$; or $\alpha\geq\max\{1/p-(n-1)/2,(1-n)/p\}$ for $2<p<\infty$. Furthermore, we improve the result of Kohen [13, Theorem 3] by showing that $\mathfrak{m}^{\alpha}$ is bounded on $L^{p}(\mathbb{H}^{n})$ provided that $\mathop{\mathrm{Re}}\alpha>\max\{{(2-n)/p}-{1/(pp_{n})},\ {(2-n)/p}-(p-2)/[pp_{n}(p_{n}-2)]\}$ for $2\leq p\leq\infty$, with $p_{n}=2(n+1)/(n-1)$ for $n\geq 3$ and $p_{n}=4$ for $n=2$.
Key words and phrases:
Spherical maximal operators, hyperbolic spaces, Fourier integral operators, Fourier multipliers, local smoothing.
2010 Mathematics Subject Classification:
43A85, 22E30, 43A90
1. Introduction
1.1. Background
The spherical maximal operator of (complex) order $\alpha$ on ${\mathbb{R}^{n}}$ was introduced in 1976 by Stein [24], which is defined by
(1.1) ${\mathfrak{M}}^{\alpha}f(x)=\sup_{t>0}|{\mathfrak{M}}^{\alpha}_{t}f(x)|,$
where ${\mathfrak{M}}^{\alpha}_{t}$ is the spherical operator given by
(1.2) ${\mathfrak{M}}^{\alpha}_{t}f(x)={1\over\Gamma(\alpha)}\int_{\{|y|\leq 1\}}\left(1-{|y|^{2}}\right)^{\alpha-1}f(x-ty)\,\mathrm{d}y.$
In the case $\alpha=1$, ${\mathfrak{M}}^{\alpha}$ corresponds to the Hardy-Littlewood maximal operator and in the case $\alpha=0$, one recovers the spherical maximal operator ${\mathfrak{M}}f(x)=\sup_{t>0}|{\mathfrak{M}}_{t}f(x)|$ in which
(1.3) ${\mathfrak{M}}_{t}f(x)={1\over\omega_{n-1}}\int_{{\mathbb{S}}^{n-1}}f(x-t\omega)\,\text{d}\omega,\ \ \ x\in\mathbb{R}^{n},\ t>0,$
[…]
(i)  In [24, Theorem 2], Stein showed that for $n\geq 2$
(1.4) $\|{\mathfrak{M}}^{\alpha}f\|_{L^{p}({\mathbb{R}^{n}})}\leq C\|f\|_{L^{p}({\mathbb{R}^{n}})}$
provided that $\alpha>1-n+n/p$ when $1<p\leq 2;$ or $\alpha>(2-n)/p$ when $2\leq p\leq\infty.$ The above maximal theorem tells us that when $\alpha=0$ and $n\geq 3$, the operator ${\mathfrak{M}}$ is bounded on $L^{p}({\mathbb{R}^{n}})$ for $p>n/(n-1)$. This range of $p$ is sharp. As is pointed out in [24, 25], no such result can hold for $p\leq n/(n-1)$ if $n\geq 2$.
(ii)  In [2], Bourgain proved that when $\alpha=0$ and $n=2,$ the operator ${\mathfrak{M}}$ is bounded on $L^{p}{(\mathbb{R}^{2})}$ for $p>2$. Bourgain’s theorem indicates that there exists $\varepsilon(p)>0$ such that
[…]
This result cannot hold even for $\alpha=0$ when $p=2$, see [24, 25]. Subsequently Mockenhaupt, Seeger and Sogge [16] gave an alternative proof of Bourgain’s result by using local smoothing estimates for the wave operator in $2+1$ dimensions.
(iii)  Using the Bourgain-Demeter decoupling theorem [3], Miao, Yang and Zheng [15] extended certain range of $\alpha$ in Stein’s result [24, Theorem 2] by showing that for $n\geq 2$ and $p>2$, (1.4) holds whenever $\mathop{\mathrm{Re}}\alpha>\max\left\{{(1-n)/4}+{(3-n)/(2p)},\,{(1-n)/p}\right\}.$
(iv)  Recently, Liu, Song, and the second and fourth named authors [14] obtained the necessary condition on $(\alpha,p)$ for the operator ${\mathfrak{M}}^{\alpha}$ to be bounded on $L^{p}(\mathbb{R}^{n})$ by showing that for $n\geq 2$ and $p>2$ (1.4) holds only if $\mathop{\mathrm{Re}}\alpha\geq\max\{1/p-(n-1)/2,\ -(n-1)/p\}.$ In the two dimensional case $n=2$, they showed that (1.4) holds whenever $\mathop{\mathrm{Re}}\alpha>\max\{1/p-1/2,\ -1/p\}$ by applying the work of Guth-Wang-Zhang [8] on local smoothing estimates for the wave operator in $2+1$ dimensions.
Note that in [17], Nowak, Roncal and Szarek [17] found sharp conditions for the operator ${\mathfrak{M}}^{\alpha}$ on radial functions to be bounded on $L^{p}_{\rm rad}({\mathbb{R}}^{n})$. They proved that for $n\geq 2$ and $\alpha>(1-n)/2$, ${\mathfrak{M}}^{\alpha}$ is bounded on $L^{p}_{\rm rad}({\mathbb{R}}^{n})$ if and only if $\alpha>1-n+n/p$ if $1<p\leq 2$; $\alpha>1/p+(1-n)/2$ if $2\leq p\leq 2n/(n-1)$; and $\alpha\geq(1-n)/p$ if $p>2n/(n-1)$.
In the last decades the spherical maximal operators have attracted a lot of attention and have been studied extensively by many authors. The maximal theorem for spherical operator has some applications; for example, Stein [24] used it to derive a Fatou’s theorem for wave equation. The spherical operator can be extended to a more general hypersurface in ${\mathbb{R}^{n}}$, see [2, 7, 10, 12, 16, 20, 21, 22, 23, 25, 26] for the references therein.
1.2. Main results
The purpose of this article is to investigate the spherical maximal operators of (complex) order $\alpha$ on the $n$-dimensional hyperbolic space $\mathbb{H}^{n}$, which was introduced and studied by Kohen ([13]). To do this, we denote the Minkowski metric on ${\mathbb{R}^{n+1}}$ by
$[z,w]=z_{0}w_{0}-z_{1}w_{1}-\cdots-z_{n}w_{n},\ \ \ \ z=(z_{0},\dots,z_{n}),w=(w_{0},\dots,w_{n})\in{\mathbb{R}^{n+1}}.$
The hyperbolic space $\mathbb{H}^{n}$ is given by
$\mathbb{H}^{n}=\{z=(z_{0},z^{\prime})\in{\mathbb{R}^{n+1}}:[z]:=[z,z]=1,\ z_{0}>0\},$
and we equip this space with the Riemannian metric induced by the above Minkowski metric [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ]. This Riemannian metric induces in turn a measure, by which we denote $\mathrm{d}z$. Define the spherical maximal operators by
(1.6) ${\mathfrak{m}}^{\alpha}(f)(z)=\sup_{t>0}|M^{\alpha}_{t}(f)(z)|,$
where
(1.7) $M^{\alpha}_{t}(f)(z)=2\mathrm{e}^{t}\left({\mathrm{e}^{t}-1\over\sinh t}\right)^{n-2}{1\over(\mathrm{e}^{t}-1)^{2\alpha+n-2}}{1\over\Gamma(\alpha)}\int_{B_{t}(z)}[\mathrm{e}^{t}z-w]^{\alpha-1}f(w)\,\mathrm{d}w,$
$B_{t}(z)$ stands for the geodesic ball in $\mathbb{H}^{n}$ centered at $z\in\mathbb{H}^{n}$ with radius $t$ and $f$ is in the Schwartz space on $\mathbb{H}^{n}$. These operators $\mathfrak{m}^{\alpha}$ and $M^{\alpha}_{t}$ are initially defined for $\mathop{\mathrm{Re}}\alpha>0,$ but the definition can be extended to all complex $\alpha$ with $\mathop{\mathrm{Re}}\alpha>(1-n)/2$ by analytic continuation, see Lemma 2.3 below. For $\alpha=0,$ one recovers the spherical maximal operator
(1.8) ${\mathfrak{m}}(f)(z)$ $=$ $\sup_{t>0}|f\ast\mathrm{d}\sigma_{t}(z)|,$
where $\mathrm{d}\sigma_{t}$ is the normalized spherical measure defined on [10, p. 277].
A natural question is for which $\alpha$ the spherical maximal opeators ${\mathfrak{m}}^{\alpha}$ are bounded on $L^{p}$. In [13, Theorem 3], Kohen showed that for $n\geq 2$,
(1.9) $\|{\mathfrak{m}}^{\alpha}f\|_{L^{p}(\mathbb{H}^{n})}\leq C\|f\|_{L^{p}(\mathbb{H}^{n})}$
holds under the following circumstances:
[…]
or
[…]
where the two dimensional case was left open since the necessary condition $p>2$ for $\alpha=0$ eliminates the use of the $L^{2}$ argument based on Plancherel’s theorem. It was not until 2000 that Ionescu [10] showed that ${\mathfrak{m}}^{0}$ is bounded on $L^{p}(\mathbb{H}^{2})$ for $2<p\leq\infty$ by making use of the local smoothing estimates for Fourier integral operators ([20]). The above two admissible relations for $\alpha$ and $p$ are summarized when $n\geq 3$ in Figure 1 below, where the relations (1.10) and (1.11) correspond to the dashed segments $AB$ and $OB$, respectively.
[…]
Figure 1. The $(1/p,\alpha)$-plot for $n\geq 3$. Kohen’s result [13, Theorem 3] indicates that $\mathfrak{m}^{\alpha}$ is bounded on $L^{p}(\mathbb{H}^{n})$ if $(1/p,\alpha)$ is strictly above the dashed folded segments $OBA$. We extend the range of $(1/p,\alpha)$ in Kohen’s result for $2<p<\infty$ to the range strictly above the dotted folded segments $OCB$. In addition, we showed that $\mathfrak{m}^{\alpha}$ is bounded on $L^{p}(\mathbb{H}^{n})$ only if $(1/p,\alpha)$ is on or above the solid folded segments $ODB$ for $2<p<\infty$, or strictly above the dashed segment $BA$ for $1<p\leq 2$.
The present article can be viewed as a continuation of the above body of work in [10, 13]. Firstly, we give a necessary condition on $(\alpha,p)$ for $\mathfrak{m}^{\alpha}$ to be bounded on $L^{p}(\mathbb{H}^{n})$ for $\alpha\in\mathbb{R}$.
Theorem 1.1.
Let $\alpha\in\mathbb{R}$ and $n\geq 2$. Suppose $\mathfrak{m}^{\alpha}$ is bounded on $L^{p}(\mathbb{H}^{n})$ for $1<p<\infty$. Then we must have:
[…]
or
[…]
Secondly, we extend the range of $\alpha$ in (1.11) for $2<p<\infty$ to the following result.
Theorem 1.2.
Suppose $2<p<\infty$ and $n\geq 2$. Then $\mathfrak{m}^{\alpha}$ is bounded on $L^{p}(\mathbb{H}^{n})$ provided that
$\mathop{\mathrm{Re}}\alpha>\max\left\{{2-n\over p}-{1\over pp_{n}},\ {2-n\over p}-{p-2\over pp_{n}(p_{n}-2)}\right\},$
where
(1.14) $p_{n}=\begin{cases}\displaystyle{2(n+1)\over n-1}&\mbox{for }n\geq 3,\\ 4&\mbox{for }n=2.\end{cases}$
Theorems  1.1 and  1.2, can be illustrated using Figure 1 above. The relations (1.12) and (1.13) in Theorem 1.1 correspond to the dashed segment $AB$ and the solid folded segments $ODB$, respectively. Theorem 1.2 indicates that compared to (1.11), there is an $1/(pp_{n})$-downwards extension of the range of $\mathop{\mathrm{Re}}\alpha$ so that (1.9) is valid for $p\geq p_{n}$, and there is a $(p-2)/[pp_{n}(p_{n}-2)]$-improvement for $2<p<p_{n}$. This is indicated in Figure 1 above by the dotted folded segments $OCB$.
The paper is organized as follows. In Section 2 we give some preliminary facts about the space $\mathbb{H}^{n}$ as well as its Fourier theory. In Section 3 we show Theorem 1.1 by constructing three examples. In Section 4 we prove Theorem 1.2 by applying an approach of Ionescu [10], whose core is to adopt the Iwasawa coordinates to use the work of Beltran-Hickman-Sogge [1] and Gao-Liu-Miao-Xi [5] on local smoothing estimates for Fourier integral operators on ${\mathbb{R}^{n}}$.
2. Fourier analysis on the hyperbolic space
In this section we recall some basic facts about the hyperbolic space $\mathbb{H}^{n}$ as well as its Fourier theory (see for examples, [9, 11, 18]). The Minkowski metric on ${\mathbb{R}^{n+1}}$ is given by
$[z,w]=z_{0}w_{0}-z_{1}w_{1}-\cdots-z_{n}w_{n},\ \ \ \ \ z=(z_{0},\dots,z_{n}),\ w=(w_{0},\dots,w_{n})\in{\mathbb{R}^{n+1}}.$
We define $\mathbb{H}^{n}$ as the hyperboloid
$\mathbb{H}^{n}=\{z=(z_{0},z^{\prime})\in{\mathbb{R}^{n+1}}:[z,z]=1,\ z_{0}>0\}$
and equip this space with the Riemannian metric induced by the Minkowski metric [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ]. This Riemannian metric induces in turn a measure, by which we denote $\mathrm{d}z$. We will further distinguish the point ${\bf 0}=(1,0,\dots,0)$.
The Laplace-Beltrami operator on $\mathbb{H}^{n}$ is denoted by $\Delta_{\mathbb{H}^{n}}$. It has spectrum $(-\infty,-(n-1)^{2}/4]$. For $\omega\in{\mathbb{S}}^{n-1}$, let
$b(\omega)=(1,\omega)\in{\mathbb{R}}^{n+1}.$
The analog of plane waves is provided by
$h_{\lambda,\omega}(z)=[z,b(\omega)]^{i\lambda-(n-1)/2},\ \ \ z\in\mathbb{H}^{n}$
(notice that $[z,b(\omega)]>0$ for $z\in\mathbb{H}^{n}$). They satisfy
$\Delta_{\mathbb{H}^{n}}h_{\lambda,\omega}(z)=-\left(\lambda^{2}+{(n-1)^{2}\over 4}\right)\,h_{\lambda,\omega}(z).$
The Helgason Fourier transform on $\mathbb{H}^{n}$ is defined as
(2.1) ${\mathscr{F}}(f)(\lambda,\omega)=\int_{\mathbb{H}^{n}}f(z)\,h_{\lambda,\omega}(z)\,\mathrm{d}z,\ \ \ \ (\lambda,\omega)\in{\mathbb{R}}_{+}\times{\mathbb{S}}^{n-1},$
while the inverse Fourier transform is then given by
(2.2) $f(z)={1\over\omega_{n-1}}\int_{0}^{\infty}\int_{\mathbb{S}^{n-1}}{\mathscr{F}}(f)(\lambda,\omega)\,{\overline{h_{\lambda,\omega}(z)}}\,|{\mathcal{H}}(\lambda)|^{-2}\,\mathrm{d}\lambda\mathrm{d}\omega,$
for the Harish-Chandra function
(2.3) ${\mathcal{H}}(\lambda)={2^{n-2}\Gamma(n/2)\over\sqrt{\uppi}}{\Gamma(\mathrm{i}\lambda)\over\Gamma((n-1)/2+\mathrm{i}\lambda)}.$
This time the following Plancherel’s theorem holds: the Fourier transform is an isometry from $L^{2}(\mathbb{H}^{n},\mathrm{d}z)$ to $L^{2}({\mathbb{R}_{+}}\times{\mathbb{S}^{n-1}},|{\mathcal{H}}(\lambda)|^{-2}\mathrm{d}\lambda{\mathrm{d}\omega\over\omega_{n-1}})$.
We need the estimates for the Harish-Chandra function ${\mathcal{H}}$, which will be frequently used throughout the paper. For its proof, we refer the reader to [10, Proposition A1(a)(b)].
Lemma 2.1.
Suppose $\lambda\in\mathbb{R}$. The Harish-Chandra function ${\mathcal{H}}(\lambda)$ satisfies $|{\mathcal{H}}(\lambda)|^{-2}=({\mathcal{H}}(\lambda)\,{\mathcal{H}}(-\lambda))^{-1}.$ The function $\lambda^{-1}({\mathcal{H}}(-\lambda))^{-1}$ belongs to $C^{\infty}(\mathbb{R})$ and
$\left|{\mathrm{d}^{k}\over\mathrm{d}\lambda^{k}}(\lambda^{-1}({\mathcal{H}}(\lambda))^{-1})\right|\leq C_{\alpha}\,(1+|\lambda|)^{(n-1)/2-1-k}.$
In analogy to ${\mathcal{H}}$, we define another function
(2.4) ${\mathcal{H}}^{\alpha}(\lambda)={2^{n-2}\Gamma(n/2)\over\sqrt{\uppi}}{\Gamma(\mathrm{i}\lambda)\over\Gamma((n-1)/2+\alpha+\mathrm{i}\lambda)},\ \ \ {\rm Re}\,\alpha>{(1-n)/2}.$
Specifically, when $\alpha=0$, ${\mathcal{H}}^{\alpha}(\lambda)$ is exactly the Harish-Chandra function ${\mathcal{H}}(\lambda)$ defined in (2.3). For the function ${\mathcal{H}}^{\alpha}$ we have the following result.
Lemma 2.2.
Let ${\rm Re}\,\alpha>{(1-n)/2}$ and $k\in{\mathbb{N}}$. Then the function $\lambda\,{\mathcal{H}}^{\alpha}(\lambda)$ belongs to $C^{\infty}(\mathbb{R})$ and
$\left|{\mathrm{d}^{k}\over\mathrm{d}\lambda^{k}}(\lambda\,{\mathcal{H}}^{\alpha}(\lambda))\right|\leq C_{\alpha}\,(1+|\lambda|)^{1-(n-1)/2-\mathop{\mathrm{Re}}\alpha-k}.$
Proof.
This is a consequence of Stirling’s formula ([27, p. 151, Example. (ii)]) ∎
In the space $\mathbb{H}^{n}$, we can adopt polar coordinates
$z=(\cosh r,\omega\sinh r),\ \ \ \ \ \ (r,\omega)\in{\mathbb{R}}_{+}\times{\mathbb{S}}^{n-1},$
[…]
$\mathrm{d}r^{2}+\sinh^{2}r\,\mathrm{d}s_{n-1}^{2},$
[…]
$\mathrm{d}z=(\sinh r)^{n-1}\,\mathrm{d}r\mathrm{d}\omega.$
The spherical function on $\mathbb{H}^{n}$ is defined by
$\varphi_{\lambda}(z)={1\over\omega_{n-1}}\int_{{\mathbb{S}^{n-1}}}h_{\lambda,\omega}(z)\,\mathrm{d}\omega.$
It depends only on the distance $r$ of $z$ to the origin $\mathbf{0}$ and can be written as
(2.5) $\varphi_{\lambda}(r)={\Gamma(n/2)\over\sqrt{\uppi}\,\Gamma((n-1)/2)\,}\int_{0}^{\uppi}(\cosh r-\cos s\,\cosh r)^{-(n-1)/2+\mathrm{i}\lambda}(\sin s)^{n-2}\,\mathrm{d}s.$
See [28].
Since radial functions on $\mathbb{H}^{n}$ depend on $r$ only, the Helgason Fourier transform (2.1) of such function $f$ can be written as
${\mathscr{F}}(f)(\lambda)=\int_{\mathbb{H}^{n}}f(z)\,\varphi_{\lambda}(z)\,\mathrm{d}z=\omega_{n-1}\int_{\mathbb{H}^{n}}f(r)\,\varphi_{\lambda}(r)\,(\sinh r)^{n-1}\,\mathrm{d}r$
and the inverse Fourier transform (2.2) is given by
$f(r)=\int_{0}^{\infty}{\mathscr{F}}(f)(\lambda)\,\varphi_{\lambda}(r)\,|{\mathcal{H}}(\lambda)|^{-2}\,\mathrm{d}\lambda.$
If $K$ is radial, then the convolution of $f$ and $K$ on $\mathbb{H}^{n}$ can be written as
$f\ast K(z)=\int_{\mathbb{H}^{n}}f(w)\,K(d(z,w))\,\mathrm{d}w,$
where $d(z,w)=\mathop{\mathrm{arcosh}}([z,w])$ denotes the geodesic distance between $z$ and $w$.
In the following, we denote $D=\sqrt{-\Delta_{\mathbb{H}^{n}}-(n-1)^{2}/4}$. For any bounded Borel function $m:[0,\infty)\to{\mathbb{C}}$, we define the radial Fourier multipliers $m(D)$ by the formula
(2.6) $m(D)(f)(z)={\mathscr{F}}^{-1}[m(\lambda)\,{\mathscr{F}}(f)(\lambda,\omega)](z).$
See [6, p. 10]. It follows from (2.6) that the radial Fourier multipliers $m(D)$ can be rewritten in the following:
$m(D)(f)(z)=f*K(z)=\int_{\mathbb{H}^{n}}f(w)\,K(d(z,w))\,\mathrm{d}w,$
where
$K(r)=\mathscr{F}^{-1}(m)(r)=\int_{0}^{\infty}m(\lambda)\,\varphi_{\lambda}(r)\,|{\mathcal{H}}(\lambda)|^{-2}\,\mathrm{d}\lambda.$
As mentioned in Section 1, the operator $M^{\alpha}_{t}(f)$ in (1.7) is initially defined for $\mathop{\mathrm{Re}}\alpha>0,$ but the definition can be extended to all complex $\alpha$ with $\mathop{\mathrm{Re}}\alpha>(1-n)/2$ by analytic continuation. To explain this, we shall use the Legendre function. Recall that in view of [4, p. 155, (1)], for $\mathop{\mathrm{Re}}(-\mu)\geq\mathop{\mathrm{Re}}\nu>-1$ and $\zeta\notin[-1,\infty)$ the Legendre function of order $\mu$ and degree $\nu$ has the following integral representation:
$P^{\mu}_{\nu}(\zeta)={2^{-\nu}(\zeta^{2}-1)^{-\mu/2}\over\Gamma(-\mu-\nu)\,\Gamma(\nu+1)}\int_{0}^{\infty}(\zeta+\cosh s)^{\mu-\nu-1}(\sinh s)^{2\nu+1}\,\mathrm{d}s.$
Then we have the following result.
Lemma 2.3.
The operator $M^{\alpha}_{t}(f)$ in (1.7) can be extended to $\alpha\in\mathbb{C}$ with $\mathop{\mathrm{Re}}\alpha>(1-n)/2$ via analytic continuation such that
$M^{\alpha}_{t}(f)(z)=m^{\alpha}_{t}(D)(f)(z),$
where
(2.7) $m^{\alpha}_{t}(\lambda)=2^{(n-2)/2+\alpha}\Gamma\!\left({n\over 2}\right)\mathrm{e}^{\alpha t}(\mathrm{e}^{t}-1)^{-2\alpha}(\sinh t)^{\alpha-(n-2)/2}P^{-\alpha-(n-2)/2}_{-1/2+\mathrm{i}\lambda}(\cosh t).$
Proof.
In [13, Theorem 1], Kohen showed that for $\alpha\in\mathbb{C}$ with $\mathop{\mathrm{Re}}\alpha>(1-n)/2$ one can write
$M^{\alpha}_{t}(f)(z)=m^{\alpha}_{t}(D)(f)(z)$
via analytic continuation, where
(2.8) $m^{\alpha}_{t}(\lambda)=2^{{n-2\over 2}+\alpha}\Gamma\!\left({n\over 2}\right)\mathrm{e}^{\alpha t}(\mathrm{e}^{t}-1)^{-2\alpha}(\sinh t)^{-(n-2)}I^{\alpha+{n-2\over 2}}P^{0}_{-{1\over 2}+\mathrm{i}\lambda}(\cosh t)$
and $I^{\gamma}$ is the fractional integral operator given by
$I^{\gamma}(f)(s)={1\over\Gamma(\gamma)}\int_{1}^{s}f(t)\,(s-t)^{\gamma-1}\,\mathrm{d}t.$
In view of [4, (5) on p. 159], we have
$I^{\alpha+(n-2)/2}P^{0}_{-1/2+\mathrm{i}\lambda}(\cosh t)=(\sinh t)^{\alpha+(n-2)/2}P^{-\alpha-(n-2)/2}_{-1/2+\mathrm{i}\lambda}(\cosh t).$
Substituting this back into (2.8), we obtain (2.7). ∎
We have the following asymptotics of the multiplier function $m^{\alpha}_{t}(\lambda)$, which will be used in the proof of Theorem 1.2.
Lemma 2.4.
Suppose $\lambda\in\mathbb{R}$.
  1. (i)
    If $t>0$, then
    $|m^{\alpha}_{t}(\lambda)|\leq C_{\alpha}\,(1+t)\,\mathrm{e}^{-(n-1)t/2}.$
  2. (ii)
    If $0<t\leq\uppi$ satisfying $t|\lambda|\geq 1$, then $m^{\alpha}_{\lambda}(t)$ can be written in the form
    (2.9) $m^{\alpha}_{t}(\lambda)=\mathrm{e}^{\mathrm{i}\lambda t}a^{\alpha,N}_{1}(\lambda,t)+\mathrm{e}^{-\mathrm{i}\lambda t}a^{\alpha,N}_{1}(-\lambda,t)+E^{\alpha,N}(\lambda,t),$
    where
    (2.10) $\begin{cases}\displaystyle|{\partial^{k}_{\lambda}}{\partial^{l}_{t}}a^{\alpha,N}_{1}(\lambda,t)|\leq C_{N}\,(|\lambda|t)^{-\mathop{\mathrm{Re}}\alpha-(n-1)/2}|\lambda|^{-k}t^{-l},\\[4.0pt] |E^{\alpha,N}(\lambda,t)|\leq C_{N}\,(|\lambda|t)^{-\mathop{\mathrm{Re}}\alpha-(n-1)/2-N-1}\end{cases}$
    for all integers $N,k,l\geq 0$, and for $\lambda,t$ in the ranges stated above.
  3. (iii)
    If $t>(\log 2)/2$, then $m^{\alpha}_{t}(\lambda)$ can be written in the form
    $m^{\alpha}_{t}(\lambda)=\mathrm{e}^{-(n-1)t/2}\left(\mathrm{e}^{\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(\lambda)\,a^{\alpha}_{2}(\lambda,t)+\mathrm{e}^{-\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(-\lambda)\,a^{\alpha}_{2}(-\lambda,t)\right),$
    where ${\mathcal{H}}^{\alpha}$ is as in (2.4) and $a^{\alpha}_{2}(\lambda,t)\in S^{0}_{\lambda}$, i.e.,
    $|\partial^{k}_{\lambda}\partial^{l}_{t}a^{\alpha}_{2}(\lambda,t)|\leq C_{k,l}(1+|\lambda|)^{-k}.$
Proof.
According to [4, p. 156, (8)],
$P_{-1/2+\mathrm{i}\lambda}^{-\alpha-(n-2)/2}(\cosh t)=C_{\alpha}\,(\sinh t)^{-\alpha-(n-2)/2}\int_{0}^{t}(\cosh t-\cosh s)^{\alpha+(n-3)/2}\cos(\lambda s)\,\mathrm{d}s.$
A direct calculation shows that for any integer $k\geq 0$
$|P_{-1/2+\mathrm{i}\lambda}^{-\alpha-(n-2)/2}(\cosh t)|\leq C_{\alpha}\begin{cases}t^{\mathop{\mathrm{Re}}\alpha+(n-2)/2}&\mbox{for }0<t\leq 1,\\ t\,\mathrm{e}^{-t/2}&\mbox{for }t>1.\end{cases}$
Substituting this back into (2.7), we obtain (i).
For part (ii) we invoke [19, (2.4.1(5))] to write that for $t\leq\uppi$,
(2.11) $(\sinh t)^{1/2}P_{-1/2+\mathrm{i}\lambda}^{-\alpha-(n-2)/2}(\cosh t)=t^{\alpha+(n-1)/2}\sum_{j=0}^{N}t^{2j}b^{\alpha}_{j}(t)\,\mathfrak{J}_{\alpha+(n-2)/2+j}(\lambda t)+{\mathfrak{R}}^{\alpha}(\lambda,t),$
where $|\mathrm{d}^{l}b^{\alpha}_{j}(t)/\mathrm{d}t^{l}|\leq C_{\alpha,l}$ for $l\geq 0$ uniformly in $j\geq 0$ ([19, (2.4.1(4))]),
$\mathfrak{J}_{m}(r)=\int_{-1}^{1}\mathrm{e}^{\mathrm{i}vr}(1-v^{2})^{m-1/2}\,\mathrm{d}v$
and by [19, (2.4.2(9))],
$|{\mathfrak{R}}^{\alpha}(\lambda,t)|\leq C_{\alpha}(|\lambda|t)^{-\mathop{\mathrm{Re}}\alpha-(n-1)/2-N-1}.$
It’s well known that for all $m\in\mathbb{C}$ with $\mathop{\mathrm{Re}}m>-1/2$ and $r$ large,
$\mathfrak{J}_{m}(r)=\mathrm{e}^{\mathrm{i}r}\psi_{m}(r)+\mathrm{e}^{-\mathrm{i}r}\psi_{m}(-r)+{\mathfrak{O}}_{m}(r),$
where $\psi_{m}(r)$ are smooth functions such that for all integer $k\geq 0$,
(2.12) $\left|{\mathrm{d}^{k}\psi_{m}(r)\over\mathrm{d}r^{k}}\right|\leq C_{m}|r|^{-m-1/2-k}$
and $|{\mathfrak{O}}_{m}(r)|\leq C_{m}|r|^{-m-1/2-N-1}$. See [10, p. 51]. Now set
(2.13) $a^{\alpha,N}_{1}(\lambda,t)=2^{{n-2\over 2}+\alpha}\Gamma\!\left({n\over 2}\right)\mathrm{e}^{\alpha t}(\mathrm{e}^{t}-1)^{-2\alpha}(\sinh t)^{\alpha-{n-1\over 2}}t^{\alpha+(n-1)/2}\sum_{j=0}^{N}t^{2j}b^{\alpha}_{j}(t)\,\psi_{\alpha+(n-2)/2+j}(\lambda t)$
and
$E^{\alpha,N}(\lambda,t)$ $=$ $2^{{n-2\over 2}+\alpha}\Gamma\!\left({n\over 2}\right)\mathrm{e}^{\alpha t}(\mathrm{e}^{t}-1)^{-2\alpha}(\sinh t)^{\alpha-(n-1)/2}$
×$\times\left({\mathfrak{R}}^{\alpha}(\lambda,t)+t^{\alpha+(n-1)/2}\sum_{j=0}^{N}t^{2j}b^{\alpha}_{j}(t)\,\mathfrak{O}_{\alpha+(n-2)/2+j}(\lambda t)\right),$
Substitute (2.11) back into (2.7), we obtain (2.9). Estimates (2.10) follow from (2.12) and the estimates on the error terms ${\mathfrak{R}}^{\alpha}$ and ${\mathfrak{O}}_{m}$.
To prove part (iii), it follows from the first equation on [19, p. 262] that
(2.14) $(\sinh t)^{1/2}P^{-\alpha-(n-2)/2}_{1/2+\mathrm{i}\lambda}(\cosh t)$ $=$ ${\mathcal{H}}^{\alpha}(\lambda)\,\mathrm{e}^{\mathrm{i}\lambda t}F\!\left(-\alpha-{n-3\over 2},\alpha+{n-1\over 2};1-\mathrm{i}\lambda;{-1\over\mathrm{e}^{2t}-1}\right)$
$+{\mathcal{H}}^{\alpha}(-\lambda)\,\mathrm{e}^{-\mathrm{i}\lambda t}F\!\left(-\alpha-{n-3\over 2},\alpha+{n-1\over 2};1+\mathrm{i}\lambda;{-1\over\mathrm{e}^{2t}-1}\right),$
where $F$ is the classical hypergeometric function, which has the integral representation
$F(a,b;c;\nu)={\Gamma(c)\over\Gamma(b)\Gamma(c-b)}\int_{0}^{1}s^{b-1}(1-s)^{c-b-1}(1-s\nu)^{-a}\,\mathrm{d}s\ \ \ \ \mbox{for }\mathop{\mathrm{Re}}c>\mathop{\mathrm{Re}}b>0$
by [4, p. 59, (10)]. For $\zeta\in\mathbb{C}$ and $k\in{\mathbb{N}}$, we denote $(\zeta)_{k}=\prod_{l=1}^{k}(\zeta+l-1)$. It follows from [4, p. 76, (10)] that for $t>(\log 2)/2$,
$F\!\left(-\alpha-{n-3\over 2},\alpha+{n-1\over 2};1+\mathrm{i}\lambda;{-1\over\mathrm{e}^{2t}-1}\right)=\sum_{k\geq 0}{(-\alpha-(n-3)/2)_{k}(\alpha+(n-1)/2)_{k}\over(1+\mathrm{i}\lambda)_{k}\,k!}\left({-1\over\mathrm{e}^{2t}-1}\right)^{k}$
belongs to $S^{0}_{\lambda}$. Now we substitute (2.14) back into (2.7) to write
$m^{\alpha}_{t}(\lambda)=\mathrm{e}^{-(n-1)t/2}\left(\mathrm{e}^{\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(\lambda)\,a^{\alpha}_{2}(\lambda,t)+\mathrm{e}^{-\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(-\lambda)\,a^{\alpha}_{2}(-\lambda,t)\right),$
where
$a^{\alpha}_{2}(\lambda,t)$ $=$ $2^{(n-2)/2+\alpha}\Gamma\!\left({n\over 2}\right)\mathrm{e}^{(\alpha+(n-1)/2)t}(\mathrm{e}^{t}-1)^{-2\alpha}(\sinh t)^{\alpha-(n-1)/2}$
×$\times F\!\left(-\alpha-{n-3\over 2},\alpha+{n-1\over 2};1-\mathrm{i}\lambda;{-1\over\mathrm{e}^{2t}-1}\right)$
belongs to $S^{0}_{\lambda}$. The proof of (iii) is concluded. ∎
Remark 2.5.
When $\alpha=0$, $m^{\alpha}_{t}(\lambda)$ is exactly a constant multiple of the spherical function $\varphi_{\lambda}(t)$. Therefore, $\varphi_{\lambda}(t)$ has the above asymptotics in Lemma 2.4 with $\alpha=0$.
3. Proof of Theorem 1.1
To prove Theorem 1.1, it reduces to show the following Lemmas 3.1, 3.2 and 3.3. To do this, we set $\rho(\cdot)=d(\cdot,\mathbf{0})$ in which $\mathbf{0}=(1,0,\cdots,0)$.
Note that for $z,w\in\mathbb{H}^{n}$
$[\mathrm{e}^{t}z-w]=2\mathrm{e}^{t}(\cosh t-\cosh d(z,w)).$
Also, for $0\leq s\leq t\leq 1$, we have
$\cosh t-\cosh s={\mathrm{e}^{t}\over 2}(1-\mathrm{e}^{s-t})\,(1-\mathrm{e}^{-s-t})\sim t^{2}-s^{2}.$
These relations will be used often in this section.
Lemma 3.1.
Suppose $\alpha\in\mathbb{R}$ and $\mathfrak{m}^{\alpha}$ is bounded on $L^{p}(\mathbb{H}^{n})$ for $1<p<\infty$. Then we must have $\alpha>1-n+n/p$.
Proof.
Let $\Gamma$ be the cone vertexed at $\mathbf{0}$ and tangent to the ball $B_{3c_{1}}(\mathbf{0}\cdot\mathfrak{a}(1/2))$, where $c_{1}$ is a constant to be determined and $\mathfrak{a}(r)$ is the Lorentz boost given by
$\mathfrak{a}(r)=\left(\begin{array}[]{ccc}\cosh r&\sinh r&0\\ \sinh r&\cosh r&0\\ 0&0&\mathrm{Id}_{n-1}\end{array}\right).$
Recall that $\rho(\cdot)=d(\cdot,\mathbf{0})$. Define $f_{\delta}$ by
$f_{\delta}(w)=\begin{cases}\displaystyle{(\rho(w))^{1-n-\alpha}\over-\log\rho(w)}&\mbox{in }(B_{1/2}(\mathbf{0})\setminus B_{\delta}(\mathbf{0}))\cap\Gamma,\\[6.0pt] 0&\mbox{outside }(B_{1/2}(\mathbf{0})\setminus B_{\delta}(\mathbf{0}))\cap\Gamma\end{cases}$
for $\delta$ sufficiently small.
Now we estimate $M^{\alpha}_{\rho(z)}(f_{\delta})(z)$ for $z\in B_{c_{1}}(\mathbf{0}\cdot\mathfrak{a}(1/2))$. To do this we let $\widetilde{S}_{z}\subset\Gamma$ be a cone vertexed at $\mathbf{0}$ and tangent to $B_{c_{1}}(z)$. Set $S_{z}=B_{1/10}(\mathbf{0})\cap\widetilde{S}_{z}$. We choose $c_{1}$ sufficiently small such that for all sufficiently small $\delta>0$ and $z\in B_{c_{1}}(\mathbf{0}\cdot a(1/2))$, $S_{z}\cap\mathop{\mathrm{supp}}f_{\delta}$ is contained in the interior of $B_{\rho(z)}(z)$, and that for $w\in S_{z}\cap\mathop{\mathrm{supp}}f_{\delta}$ we have
$\rho(z)-d(z,w)\sim\rho(w)>0.$
Thus in view of (1.7) for such $z$ we can write
$|M^{\alpha}_{\rho(z)}(f_{\delta})(z)|$$\geq$ $C_{\alpha}\int_{S_{z}\cap\mathop{\mathrm{supp}}f_{\delta}}(\cosh\rho(z)-\cosh d(z,w))^{\alpha-1}{(\rho(w))^{1-n-\alpha}\over-\log(\rho(w))}\,\mathrm{d}w$
$\geq$ $C_{\alpha}\int_{\delta}^{1/10}r^{\alpha-1}{r^{1-n-\alpha}\over-\log r}r^{n-1}\,\mathrm{d}r$
$\geq$ $C_{\alpha}\log\log(1/\delta)$
using polar coordinates. Hence
$\|\mathfrak{m}^{\alpha}(f_{\delta})\|_{p}\geq\left(\int_{B_{c_{1}}(\mathbf{0}\cdot\mathfrak{a}(1/2))}|M^{\alpha}_{\rho(z)}(f_{\delta})(z)|^{p}\,\mathrm{d}z\right)^{1/p}\geq C_{\alpha}\log\log(1/\delta).$
On the other hand,
$\|f_{\delta}\|_{p}$$\leq$ $C\left(\int_{\delta}^{1/2}{r^{(1-n-\alpha)p}\over(-\log r)^{p}}r^{n-1}\,\mathrm{d}r\right)^{1/p}$
$\leq$ $\begin{cases}C_{\alpha}&\mbox{for }\alpha\leq 1-n+n/p,\\[4.0pt] C_{\alpha}\delta^{1-n-\alpha+n/p}&\mbox{for }\alpha>1-n+n/p.\end{cases}$
If $\|\mathfrak{m}^{\alpha}(f)\|_{p}\leq C\|f\|_{p}$ for all $f$, then we must have $\alpha>1-n+n/p$. ∎
Lemma 3.2.
Suppose $\alpha\in\mathbb{R}$ and $\mathfrak{m}^{\alpha}$ is bounded on $L^{p}(\mathbb{H}^{n})$ for $1<p<\infty$. Then we must have $\alpha\geq 1/p-(n-1)/2$.
Proof.
For $j\in{\mathbb{N}}$ sufficiently large, we set $R_{j}=\{(w_{0},w_{1},w^{\prime\prime})\in\mathbb{H}^{n}:|w_{1}-2\cdot 2^{-j}|\leq 2^{-j},|w^{\prime\prime}|\leq c_{2}2^{-j/2}\}$, where $c_{2}$ is a constant to be determined. We define $g_{j}(w)=\mathbbm{1}_{R_{j}}(w)$. Recall that $\rho(\cdot)=d(\cdot,\mathbf{0})$.
Now we consider $z=(z_{0},z_{1},z^{\prime\prime})\in\mathbb{H}^{n}$ with $\rho(z)\sim 1$ and $|z^{\prime\prime}|\leq c_{2}2^{-j/2}$. We choose $c_{2}$ sufficiently small such that for all sufficiently large integer $j$ and $z$ as above, $\mathop{\mathrm{supp}}g_{j}$ is contained in the interior of $B_{\rho(z)}(z)$. Under this circumstance, if $w\in\mathop{\mathrm{supp}}g_{j}$, then we have $\rho(z)-d(z,w)>0$. Thus in view of (1.7) for such $z$ we can write
$|M^{\alpha}_{\rho(z)}(g_{j})(z)|$$\geq$ $C_{\alpha}\int_{\mathop{\mathrm{supp}}g_{j}}(\cosh\rho(z)-\cosh d(z,w))^{\alpha-1}\,\mathrm{d}w$
$\geq$ $C_{\alpha}\int_{\rho(z)-c2^{-j}}^{\rho(z)-c^{\prime}2^{-j}}2^{-(n-1)j/2}(\rho(z)-r)^{\alpha-1}\,\mathrm{d}r$
$\geq$ $C_{\alpha}2^{-(\alpha+(n-1)/2)j}$
using polar coordinates, with $c,c^{\prime}>0$ independent of $j$. Hence
$\|\mathfrak{m}^{\alpha}(g_{j})\|_{p}$$\geq$ ($\left(\int_{\{z=(z_{0},z_{1},z^{\prime\prime})\in\mathbb{H}^{n}:\rho(z)\sim 1,|z^{\prime\prime}|\leq c_{2}2^{-j/2}\}}|M^{\alpha}_{\rho(z)}(g_{j})(z)|^{p}\,\mathrm{d}z\right)^{1/p}$
$\geq$ $C_{\alpha}2^{-(n-1)j/(2p)}2^{-(\alpha+(n-1)/2)j}.$
On the other hand, $\|g_{j}\|_{p}\leq C2^{-(n+1)j/(2p)}$. If $\|\mathfrak{m}^{\alpha}(f)\|_{p}\leq C\|f\|_{p}$ for all $f$, then we must have
$2^{-(n-1)j/(2p)}2^{-(\alpha+(n-1)/2)j}\leq 2^{-(n+1)j/(2p)},$
which implies exactly $\alpha\geq 1/p+(1-n)/2$. ∎
Lemma 3.3.
Suppose $\alpha\in\mathbb{R}$ and $\mathfrak{m}^{\alpha}$ is bounded on $L^{p}(\mathbb{H}^{n})$ for $1<p<\infty$. Then we must have $\alpha\geq(1-n)/p$.
Proof.
Recall that $\rho(\cdot)=d(\cdot,\mathbf{0})$. Set $E_{\varepsilon}=\{w\in\mathbb{H}^{n}:|\rho(w)-1+2\varepsilon|<\varepsilon\}$ for $\varepsilon$ sufficiently small. We define $h_{\varepsilon}(w)=\mathbbm{1}_{E_{\varepsilon}}(w)$.
Now we consider $z\in B_{\varepsilon/100}(\mathbf{0})$. Then it’s obvious that $\mathop{\mathrm{supp}}h_{\varepsilon}$ is contained in the interior of $B_{1}(z)$, which indicates that if $w\in\mathop{\mathrm{supp}}h_{\varepsilon}$, then we have $01-d(z,w)>$ 0. Thus in view of (1.7) for such $z$ we can write
$|M^{\alpha}_{1}(h_{\varepsilon})(z)|$$\geq$ $C_{\alpha}\int_{\mathop{\mathrm{supp}}h_{\varepsilon}}(\cosh 1-\cosh d(z,w))^{\alpha-1}\,\mathrm{d}w$
$\geq$ $C_{\alpha}\int_{1-\widetilde{c}\varepsilon}^{1-\widetilde{c}^{\prime}\varepsilon}(1-r)^{\alpha-1}\,\mathrm{d}r\geq C_{\alpha}\varepsilon^{\alpha}$
using polar coordinates, with $\widetilde{c},\widetilde{c}^{\prime}>0$ independent of $\varepsilon$. Hence
$\|\mathfrak{m}^{\alpha}(h_{\varepsilon})\|_{p}\geq\left(\int_{B_{\varepsilon/100}(\mathbf{0})}|M^{\alpha}_{1}(h_{\varepsilon})(z)|^{p}\,\mathrm{d}z\right)^{1/p}\geq C_{\alpha}\varepsilon^{n/p}\varepsilon^{\alpha}.$
On the other hand, $\|h_{\varepsilon}\|_{p}\leq C\varepsilon^{1/p}$. If $\|\mathfrak{m}^{\alpha}(f)\|_{p}\leq C\|f\|_{p}$ for all $f$, then we must have
$\varepsilon^{n/p}\varepsilon^{\alpha}\leq\varepsilon^{1/p},$
which is exactly $\alpha\geq(1-n)/p$. ∎
4. Proof of Theorem 1.2
To begin with, we state some facts about the local smoothing estimates for Fourier integral operators ([21, Chapter 8]), which will be used in the proof of Theorem 1.2. To do this, we let $\psi_{0},\psi_{1}\in C_{c}^{\infty}(\mathbb{R}^{n})$ having sufficiently small supports, and suppose $\Phi(x,y)$ is a smooth function defined on $\mathop{\mathrm{supp}}\psi_{0}\times\mathop{\mathrm{supp}}\psi_{1}$ such that the Monge-Ampere matrix associated to $\Phi$ is non-singular:
(4.1) $\left|\det\!\left(\begin{array}[]{cc}0&\partial\Phi/\partial x\\ \displaystyle{\partial\Phi\over\partial y}&\displaystyle{\partial^{2}\Phi\over\partial x\partial y}\end{array}\right)\right|\geq c>0,\ \ \ \ \ x\in\mathop{\mathrm{supp}}\psi_{0},\ y\in\mathop{\mathrm{supp}}\psi_{1},$
and there is a homogeneous function $q(x,\cdot)$ with $\mathop{\mathrm{rank}}\mathop{\mathrm{Hess}}q(x,\cdot)\,\equiv n-1$ such that
(4.2) $q\big{(}x,\Phi_{x}^{{}^{\prime}}(x,y)\big{)}\,\equiv 1,\ \ \ \ \ x\in\mathop{\mathrm{supp}}\psi_{0},\ y\in\mathop{\mathrm{supp}}\psi_{1}.$
The property (4.1) is usually referred to as rotational curvature while the property (4.2) is a simplified version of Sogge’s cinematic curvature hypothesis. It follows from [20, p. 354] that the above $\Phi(x,y)$ in (4.2) satisfies a Carleson-Sjölin type hypothesis.
Given a symbol $a(\lambda,t,x,y)\in S^{m}_{\lambda}$, i.e., $a\in C^{\infty}(\mathbb{R}\times\mathbb{R}\times\mathbb{R}^{n}\times\mathbb{R}^{n})$ satisfying
(4.3) $|\partial^{k}_{\lambda}\partial^{l}_{t}\partial^{\gamma}_{x,y}a(\lambda,t,x,y)|\leq C_{k,\gamma,l}\,(1+|\lambda|)^{m-k}\qquad\mbox{for }x\in\mathop{\mathrm{supp}}\psi_{0},\ y\in\mathop{\mathrm{supp}}\psi_{1}\mbox{ and }t\in I$
for some closed interval $I\subset\mathbb{R}$, we can define the associated Fourier integral operator by
$F_{t}(f)(x)=\int_{\mathbb{R}^{n}}f(y)\,K_{t}(x,y)\,\mathrm{d}y,$
where
(4.4) $K_{t}(x,y)=\psi_{0}(x)\,\psi_{1}(y)\int_{\mathbb{R}}\mathrm{e}^{\mathrm{i}\lambda(t-\Phi(x,y))}a(\lambda,t,x,y)\,\mathrm{d}\lambda.$
The following result concerning the local smoothing estimates for Fourier integral operators is a combination of [1, Proposition 3.2] and [5, Theorem 1.4]. Recall that $p_{n}$ is given in (1.14), i.e., $p_{n}={2(n+1)/(n-1)}$ for $n\geq 3$; $p_{n}=4$ for $n=2$.
Lemma 4.1.
Suppose $\Phi$ satisfies (4.1), (4.2) and $a(\lambda,t,x,y)$ satisfies (4.3) with $t\in I$ for some closed interval $I\subset\mathbb{R}$. Then for any $s<1/p_{n}$ with $p_{n}$ as in (1.14)
(4.5) $\|F_{t}(f)(x)\|_{L^{p_{n}}_{x,t}(\mathbb{R}^{n}\times I)}\leq C_{m,s}\|f\|_{W^{m-(n-1)/p_{n}-s,p_{n}}(\mathbb{R}^{n})}.$
Remark 4.2.
Beltran et al. [1] showed that for $n\geq 2$ and $2(n+1)/(n-1)\leq p<\infty$,
(4.6) $\|F_{t}(f)(x)\|_{L^{p}_{x,t}(\mathbb{R}^{n}\times I)}\leq C_{m,s,p}\|f\|_{W^{m-(n-1)/p-s,p}(\mathbb{R}^{n})},\ \ \ \ s<1/p$
by establishing a variable coefficient version of the Wolff-type decoupling estimate of Bourgain-Demeter [3]. Later, Gao et al. [5] improved this result for the case $n=2$, and they showed that (4.6) holds for $4\leq p<\infty$ and $n=2$ by establishing a variable coefficient version of the square function estimate of Guth-Wang-Zhang (see [8]).
Let $\beta\in C_{c}^{\infty}(\mathbb{R})$ with $\mathop{\mathrm{supp}}\beta\subset[-2,-1/2]\cup[1/2,2]$ such that $\sum_{j\in\mathbb{Z}}\beta(2^{-j}\cdot)\,\equiv 1$ on $\mathbb{R}\setminus\{0\}$. We set $\beta_{j}=\beta(2^{-j}\cdot)$ for all integrals $j>0$ and $\beta_{0}=1-\sum_{j\geq 1}\beta_{j}$. Define
$F_{t,j}(f)(x)=\int_{\mathbb{R}^{n}}f(y)\,K_{t,j}(x,y)\,\mathrm{d}y,$
where
$K_{t,j}(x,y)=\psi_{0}(x)\,\psi_{1}(y)\int_{\mathbb{R}}\beta_{j}(\lambda)\,\mathrm{e}^{\mathrm{i}\lambda(t-\Phi(x,y))}a(\lambda,t,x,y)\,\mathrm{d}\lambda.$
The following Proposition 4.3 can be obtained by minor modifications of the proof of [21, Theorems 6.3.1 and 8.3.1], so we omit it here.
Proposition 4.3.
Suppose $\Phi$ satisfies (4.1), (4.2) and $a(\lambda,t,x,y)$ satisfies (4.3) with $t\in I$ for some closed interval $I\subset\mathbb{R}$. Then for every sufficiently small $\delta>0$ we have
(4.7) $\left\|\sup_{t\in I}|F_{t,j}(f)|\right\|_{p_{n}}\leq C_{m,\delta}2^{(m-(n-1)/p_{n}+\delta)j}\|f\|_{p_{n}}.$
As a consequence, we have
(4.8) $\left\|\sup_{t\in I}|F_{t}(f)|\right\|_{p_{n}}\leq C_{m}\|f\|_{p_{n}}$
provided that $m<(n-1)/p_{n}$.
Now we apply Proposition 4.3 to prove the following result, which implies Theorem 1.2 as a special case with $p=p_{n}$. Precisely, we have
Theorem 4.4.
Let $p_{n}$ be as in (1.14). Then $\mathfrak{m}^{\alpha}$ is bounded on $L^{p_{n}}(\mathbb{H}^{n})$ whenever $\mathop{\mathrm{Re}}\alpha>(2-n)/p_{n}-1/p_{n}^{2}$.
To prove Theorem 4.4, it is sufficient to show the following Proposition 4.5 for the “global” part of the operator $\mathfrak{m}^{\alpha}$ and Proposition 4.10 for the “local” part. Roughly speaking, our approach in the proof of Proposition 4.5 is inspired by an argument of Ionescu in [10], in which the crucial step is to adopt the Iwasawa coordinates to use the local smoothing of Fourier integral operators. We will prove Proposition 4.10 by using the local smoothing of Fourier integral operators.
Proposition 4.5.
Let $p_{n}$ be as in (1.14). Then for every $T\geq 10$ and every small $\delta>0$,
(4.9) $\left\|\sup_{t\in[T,T+1]}|M^{\alpha}_{T}(f)|\right\|_{p_{n}}\leq C_{\alpha,\delta}\mathrm{e}^{-(\mathop{\mathrm{Re}}\alpha+(n-2)/p_{n}+1/p_{n}^{2}-\delta)(n-1)p_{n}T}\|f\|_{p_{n}}.$
As a consequence, we have
(4.10) $\left\|\sup_{t>10}|M^{\alpha}_{t}(f)|\right\|_{p_{n}}\leq C_{\alpha}\|f\|_{p_{n}}$
provided that $\mathop{\mathrm{Re}}\alpha>(2-n)/p_{n}-1/p_{n}^{2}$.
Let us see how to prove Proposition 4.5. Assume $T\geq 10$ and set
$\mathfrak{m}^{\alpha}_{T}(f)(z)=\sup_{t\in[T,T+1]}|M^{\alpha}_{t}(f)(z)|.$
It follows from (iii) of Lemma 2.4 that for $t\in[T,T+1]$,
$m^{\alpha}_{t}(\lambda)=\mathrm{e}^{-(n-1)t/2}\left(\mathrm{e}^{\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(\lambda)\,a^{\alpha}_{2}(\lambda,t)+\mathrm{e}^{-\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(-\lambda)\,a^{\alpha}_{2}(-\lambda,t)\right),$
where $a^{\alpha}_{2}(\lambda,t)\in S^{0}_{\lambda}$ and ${\mathcal{H}}^{\alpha}$ is as in (2.4). Set $\phi_{s}\in C^{\infty}_{c}(\mathbb{R})$, $0\leq\phi_{s}\leq 1$, $\phi_{s}\equiv 1$ on $[0,s+1]$ and $\phi_{s}\equiv 0$ outside $(-1,s+2)$. The convolution kernel $K^{\alpha}_{t}(r)$ of $M^{\alpha}_{t}$ can be written as the following:
$K^{\alpha}_{t}(r)=\mathscr{F}^{-1}(m^{\alpha}_{t})(r)=\phi_{T}(r)\int_{0}^{\infty}m^{\alpha}_{t}(\lambda)\,\varphi_{\lambda}(r)\,|{\mathcal{H}}(\lambda)|^{-2}\,\mathrm{d}\lambda$
where we used the fact that $\mathop{\mathrm{supp}}K^{\alpha}_{t}(r)\subset\{r:0\leq r\leq t\}$, see (1.7). Define
(4.11) $K^{\alpha}_{t,j}(r)=\phi_{T}(r)\int_{0}^{\infty}\beta_{j}(\lambda)\,m^{\alpha}_{t}(\lambda)\,\varphi_{\lambda}(r)\,|{\mathcal{H}}(\lambda)|^{-2}\,\mathrm{d}\lambda$
so that $K^{\alpha}_{t,j}(r)=\sum\limits_{j\geq 0}K^{\alpha}_{t,j}(r)$. By Lemma 2.3 and the asymptotics of $m^{\alpha}_{t}$ and $\varphi_{\lambda}$ (see Lemma 2.4(iii) and Remark 2.5), we rewrite $K^{\alpha}_{t,j}(r)=\mathscr{A}^{\alpha}_{t,j}(r)+\mathscr{E}^{\alpha}_{t,j}(r),$ where
(4.12) $\mathscr{A}^{\alpha}_{t,j}(r)$ $=$ $(\phi_{T}(r)-\phi_{0}(r))\,\mathrm{e}^{-(n-1)(r+t)/2}$
×$\times\int_{\mathbb{R}}\beta_{j}(\lambda)\,{\mathcal{H}}^{\alpha}(\lambda)\left({\mathrm{e}^{\mathrm{i}\lambda(t+r)}\over{\mathcal{H}}(-\lambda)}a_{2}^{0}(\lambda,r)+{\mathrm{e}^{\mathrm{i}\lambda(t-r)}\over{\mathcal{H}}(\lambda)}a_{2}^{0}(-\lambda,r)\right)a_{2}^{\alpha}(\lambda,t)\,\mathrm{d}\lambda$
and
(4.13) $\mathscr{E}^{\alpha}_{t,j}(r)$ $=$ $\phi_{0}(r)\,\mathrm{e}^{-(n-1)t/2}$
×$\times\int_{0}^{\infty}\beta_{j}(\lambda)\left(\mathrm{e}^{\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(\lambda)\,a^{\alpha}_{2}(\lambda,t)+\mathrm{e}^{-\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(-\lambda)\,a^{\alpha}_{2}(-\lambda,t)\right)\varphi_{\lambda}(r)\,|{\mathcal{H}}(\lambda)|^{-2}\,\mathrm{d}\lambda.$
Hence,
$\mathfrak{m}^{\alpha}_{T,j}(f)(z)=\sup_{t\in[T,T+1]}|f*K^{\alpha}_{t,j}(z)|\leq{\mathfrak{A}}^{\alpha}_{T,j}(f)(z)+\mathfrak{E}^{\alpha}_{T,j}(f)(z),$
where
${\mathfrak{A}}^{\alpha}_{T,j}(f)(z)$ $=$ $\sup_{t\in[T,T+1]}|f*\mathscr{A}^{\alpha}_{t,j}(z)|,$
$\mathfrak{E}^{\alpha}_{T,j}(f)(z)$ $=$ $\sup_{t\in[T,T+1]}|f*\mathscr{E}^{\alpha}_{t,j}(z)|.$
We begin with the following lemma for the error term $\mathfrak{E}^{\alpha}_{T,j}$.
Lemma 4.6.
For $1\leq p\leq\infty$, $20\leq\sigma<(n-1)/$ 2 and $j\geq 0$
(4.14) $\|\mathfrak{E}^{\alpha}_{T,j}(f)\|_{p}\leq C_{\alpha,\sigma}2^{-(\mathop{\mathrm{Re}}\alpha+\sigma)j}\mathrm{e}^{-(n-1)t/2}\|f\|_{p}.$
Proof.
By (4.13) and Lemmas 2.2 and 2.4(i)(iii),
$\sup_{t\in[T,T+1]}|\mathscr{E}^{\alpha}_{t,0}(r)|\leq C_{\alpha}\phi_{0}(r)\,\mathrm{e}^{-(n-1)t/2}.$
It follows from Schur’s lemma that (4.14) holds for the case $j=0$.
Now we assume $j\geq 1$. We take an even cutoff function $\gamma\in C^{\infty}_{c}(\mathbb{R})$ with $\gamma\equiv 1$ on $[-1,1]$ and $\gamma\equiv 0$ outside $(-2,2)$. By Lemma 2.4(ii), we write
$\mathscr{E}^{\alpha}_{t,j}(r)=\mathscr{E}^{\alpha,1}_{t,j}(r)+\mathscr{E}^{\alpha,2}_{t,j}(r)+\mathscr{E}^{\alpha,3}_{t,j}(r)$
where
$\mathscr{E}^{\alpha,1}_{t,j}(r)$ $=$ $\ \phi_{0}(r)\,\mathrm{e}^{-(n-1)t/2}$
×$\times\int_{0}^{\infty}\beta_{j}(\lambda)\,\gamma(\lambda r)\left(\mathrm{e}^{\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(\lambda)\,a^{\alpha}_{2}(\lambda,t)+\mathrm{e}^{-\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(-\lambda)\,a^{\alpha}_{2}(-\lambda,t)\right)\varphi_{\lambda}(r)\,|{\mathcal{H}}(\lambda)|^{-2}\,\mathrm{d}\lambda,$
$\mathscr{E}^{\alpha,2}_{t,j}(r)$ $=$ $\phi_{0}(r)\,\mathrm{e}^{-(n-1)t/2}$
×$\times\int_{\mathbb{R}}\beta_{j}(\lambda)\,(1-\gamma(\lambda r))\left(\mathrm{e}^{\mathrm{i}\lambda(t+r)}a^{0,N}_{1}(\lambda,r)+\mathrm{e}^{\mathrm{i}\lambda(t-r)}a^{0,N}_{1}(-\lambda,r)\right){\mathcal{H}}^{\alpha}(\lambda)\,a^{\alpha}_{2}(\lambda,t)\,|{\mathcal{H}}(\lambda)|^{-2}\,\mathrm{d}\lambda,$
$\mathscr{E}^{\alpha,3}_{t,j}(r)$ $=$ $\phi_{0}(r)\,\mathrm{e}^{-(n-1)t/2}$
×$\times\int_{0}^{\infty}\beta_{j}(\lambda)\,(1-\gamma(\lambda r))\left(\mathrm{e}^{\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(\lambda)\,a^{\alpha}_{2}(\lambda,t)+\mathrm{e}^{-\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(-\lambda)\,a^{\alpha}_{2}(-\lambda,t)\right)E^{\alpha,N}(\lambda,r)\,|{\mathcal{H}}(\lambda)|^{-2}\,\mathrm{d}\lambda.$
By Lemmas 2.1, 2.2 and 2.4 and integration by parts, taking $N$ sufficiently large, we obtain that $\mathscr{E}^{\alpha,i}_{t,j}(r)\ (i=1,2,3)$ are all bounded by a constant multiple of $\phi_{0}(r)\,\mathrm{e}^{-(n-1)t/2}2^{-(\mathop{\mathrm{Re}}\alpha+\sigma)j}r^{-\sigma-(n+1)/2}$. Hence, we obtain
(4.15) $\sup_{t\in[T,T+1]}|\mathscr{E}^{\alpha}_{t,j}(r)|\leq C_{\alpha}\phi_{0}(r)\,\mathrm{e}^{-(n-1)t/2}2^{-(\mathop{\mathrm{Re}}\alpha+\sigma)j}r^{-\sigma-(n+1)/2}.$
Note that $\mathop{\mathrm{Re}}\alpha>(1-n)/2$. By using polar coordinates, one can derive
$\sup_{z\in\mathbb{H}^{n}}\left\|\sup_{t\in[T,T+1]}|\mathscr{E}^{\alpha}_{t,j}(d(z,\cdot))|\right\|_{1}+\sup_{w\in\mathbb{H}^{n}}\left\|\sup_{t\in[T,T+1]}|\mathscr{E}^{\alpha}_{t,j}(d(\cdot,w))|\right\|_{1}\leq C_{\alpha,\sigma}\mathrm{e}^{-(n-1)t/2}$
provided that $\sigma<(n-1)/2$. Then we apply Schur’s lemma to derive the desired result (4.14). The proof of Lemma 4.6 is concluded. ∎
Next we establish $L^{p}$ boundedness of the operator $\mathfrak{m}^{\alpha}_{T,j}$ by the following three Lemmas 4.7, 4.8 and 4.9 as three special cases $p=\infty,2$ and $p_{n}$, respectively. Firstly, we have
Lemma 4.7.
For $T\geq 10$ and $j\geq 0$ we have
$\|\mathfrak{m}^{\alpha}_{T,j}(f)\|_{\infty}\leq C_{\alpha}2^{-\mathop{\mathrm{Re}}\alpha j}\|f\|_{\infty},$
Proof.
In view of (4.12), by Lemmas 2.2 and 2.4(iii) and integration by parts, for $t\in[T,T+1]$ we have
$|\mathscr{A}^{\alpha}_{t,j}(r)|\leq C_{\alpha,N}\phi_{T}(r)\,\mathrm{e}^{-(n-1)r}2^{(1-\mathop{\mathrm{Re}}\alpha)j}(1+2^{j}|t-r|)^{-N}.$
Using polar coordinates and taking $N$ sufficiently large, for $t$ as above, we have
$\sup_{w\in\mathbb{H}^{n}}\|\mathscr{A}^{\alpha}_{t,j}(d(\cdot,w))\|_{1}\leq C_{\alpha}2^{-\mathop{\mathrm{Re}}\alpha j}.$
By Hölder’s inequality,
$\|{\mathfrak{A}}^{\alpha}_{T,j}(f)\|_{\infty}\leq\|f\|_{\infty}\sup_{w\in\mathbb{H}^{n}}\sup_{t\in[T,T+1]}\|\mathscr{A}^{\alpha}_{t,j}(\cdot,w)\|_{1}\leq C_{\alpha}2^{-\mathop{\mathrm{Re}}\alpha j}\|f\|_{\infty}.$
This, in combination with Lemma 4.6 for $\sigma=0$, concludes the proof of Lemma 4.7. ∎
Lemma 4.8.
For $T\geq 10$ and $j\geq 0$ we have
$\|\mathfrak{m}^{\alpha}_{T,j}(f)\|_{2}\leq C_{\alpha}2^{-(\mathop{\mathrm{Re}}\alpha+(n-2)/2)j}\mathrm{e}^{-(n-1)T/2}(1+T)\,\|f\|_{2}.$
Proof.
By [21, Lemma 2.4.2], we see that Lemma 4.8 is a standard consequence of the following two estimates:
(4.16) $\|f*K^{\alpha}_{t,j}\|_{2}\leq C_{\alpha}2^{-(\mathop{\mathrm{Re}}\alpha+(n-1)/2)j}\mathrm{e}^{-(n-1)t/2}(1+T)\,\|f\|_{2}$
and
(4.17) $\|\partial_{t}(f*K^{\alpha}_{t,j})\|_{2}\leq C_{\alpha}2^{-(\mathop{\mathrm{Re}}\alpha+(n-3)/2)j}\mathrm{e}^{-(n-1)t/2}(1+T)\,\|f\|_{2}$
for any $t\in[T,T+1]$, where $K^{\alpha}_{t,j}$ is defined in (4.11).
Let us first verify (4.16). To do this, we set
$\widetilde{K}^{\alpha}_{t,j}(r)=(1-\phi_{T}(r))\,\mathscr{F}^{-1}(\beta_{j}m^{\alpha}_{t})(r)$
to be the complementary kernel of $K^{\alpha}_{t,j}$ such that $K^{\alpha}_{t,j}+\widetilde{K}^{\alpha}_{t,j}=\mathscr{F}^{-1}(\beta_{j}\,m^{\alpha}_{t})$. It follows from Lemmas 2.2, 2.4(i)(iii) and Plancherel’s theorem that
(4.18) $\|f*(K^{\alpha}_{t,j}+\widetilde{K}^{\alpha}_{t,j})\|_{2}\leq\|\beta_{j}m^{\alpha}_{t}\|_{\infty}\|f\|_{2}\leq C_{\alpha}2^{-(\mathop{\mathrm{Re}}\alpha+(n-1)/2)j}(1+t)\,\mathrm{e}^{-(n-1)t/2}\|f\|_{2}.$
On the other hand, we apply Lemmas 2.1, 2.4(iii) to write
$\widetilde{K}^{\alpha}_{t,j}(r)=(1-\phi_{T}(r))\,\mathrm{e}^{-(n-1)(r+t)/2}\int_{\mathbb{R}}\beta_{j}(\lambda)\left({\mathrm{e}^{\mathrm{i}\lambda(t+r)}\over{\mathcal{H}}(-\lambda)}a_{2}^{0}(\lambda,r)+{\mathrm{e}^{\mathrm{i}\lambda(t-r)}\over{\mathcal{H}}(\lambda)}a_{2}^{0}(-\lambda,r)\right)a_{2}^{\alpha}(\lambda,t)\,\mathrm{d}\lambda.$
This, together with Lemma 2.4(i) and integration by parts, implies
$\|\mathscr{F}(\widetilde{K}^{\alpha}_{t,j})\|_{\infty}\leq C_{\alpha,N}2^{-Nj}(1+t)\,\mathrm{e}^{-(n-1)t/2}$
provided $N$ sufficiently large. By Plancherel’s theorem again, we see that
(4.19) $\|f*\widetilde{K}^{\alpha}_{t,j}\|_{2}\leq\|\mathscr{F}(\widetilde{K}^{\alpha}_{t,j})\|_{\infty}\|f\|_{2}\leq C_{\alpha,N}2^{-Nj}(1+t)\,\mathrm{e}^{-(n-1)t/2}\|f\|_{2}.$
From these estimates (4.18) and (4.19), the desired estimate (4.16) follows readily. The proof of (4.17) follows from a similar argument showing (4.16). Hence the proof of Lemma 4.8 is complete. ∎
The following lemma is the essential step in proving Proposition 4.5 as a special case $p=p_{n}.$
Lemma 4.9.
Let $p_{n}$ be as in (1.14). For $T\geq 10$, $j\geq 0$ and sufficiently small $\delta>0$,
(4.20) $\left\|\mathfrak{m}^{\alpha}_{T,j}(f)\right\|_{p_{n}}\leq C_{\alpha,\delta}2^{-(\mathop{\mathrm{Re}}\alpha+(n-1)/p_{n}-\delta)j}\mathrm{e}^{(n-1)(1-1/p_{n})T}\|f\|_{p_{n}}.$
Proof.
By Lemma 4.6 with $\sigma=(n-1)/p_{n}-\delta$, it suffices to show that
(4.21) $\|{\mathfrak{A}}^{\alpha}_{T,j}(f)\|_{p_{n}}\leq C_{\alpha,\delta}2^{-\varepsilon j}\mathrm{e}^{(n-1)(1-1/p_{n})T}\|f\|_{p_{n}}$
with $\varepsilon=\mathop{\mathrm{Re}}\alpha+(n-1)/p_{n}-\delta$.
We now prove (4.21). Roughly speaking, the favorable factor $2^{-\varepsilon j}$ in (4.21) comes from the local smoothing estimate (4.7), while the unfavorable factor $\mathrm{e}^{(n-1)(1-1/p_{n})T}$ is due to several localizations we have to make and to quantitative estimates on the rotational curvature (the term on the RHS of (4.1)) of defining functions of circles of radius $\approx T$. We start by localizing the operator ${\mathfrak{A}}^{\alpha}_{T,j}$. Note that it suffices to prove that for any $\psi_{0},\psi_{1}\in C_{c}^{\infty}(\mathbb{H}^{n})$ with supports inside balls of radius $c_{0}$ ($c_{0}$ is sufficiently small to be determined later), one has
(4.22) $\|\psi_{0}\,{\mathfrak{A}}^{\alpha}_{T,j}(\psi_{1}f)\|_{p_{n}}\leq C_{\alpha,\delta}2^{-\varepsilon j}\mathrm{e}^{-(n-1)T/p_{n}}\|\mathbbm{1}_{\mathop{\mathrm{supp}}\psi_{1}}f\|_{p_{n}}.$
To see that (4.22) implies the desired estimate (4.21) we fix a partition of unity $\{\eta_{k}\}_{k\in K}$ on $\mathbb{H}^{n}$ with each $\mathop{\mathrm{supp}}\eta_{k}$ contained in a ball $B_{k}$ of radius $c_{0}$ satisfying that $(1/5)B_{k}$ are mutually disjoint. By $cB_{k}$ we mean the ball with the same center as $B_{k}$’s and radius $c$ times as $B_{k}$’s. Hence it’s obvious that
(4.23) $\left\|\sum_{k\in K}\mathbbm{1}_{2B_{k}}\right\|_{\infty}<\infty.$
For any $k\in K$ we set $L_{k}=\{l\in K:d(\mathop{\mathrm{supp}}\eta_{k},\mathop{\mathrm{supp}}\eta_{l})\leq T+2\}$. Note that
$\sup_{k\in K}\#L_{k}\leq C{\mathrm{e}^{(n-1)T}\over c_{0}^{n}}.$
Recall also that the Schwartz kernels $\mathscr{A}^{\alpha}_{t,j}(d(z,w))$ vanish unless $d(z,w)\leq T+2$. Therefore
$\|{\mathfrak{A}}^{\alpha}_{T,j}(f)\|^{p_{n}}_{p_{n}}$ $=$ $\int_{\mathbb{H}^{n}}\left|\sum_{k\in K}\eta_{k}(z)\,{\mathfrak{A}}^{\alpha}_{T,j}\!\left(\sum_{l\in L_{k}}\eta_{l}f\right)\!(z)\right|^{p_{n}}\mathrm{d}z$
$\leq$ $C\,(\#L_{k})^{p_{n}/p_{n}^{\prime}}\sum_{k\in K}\sum_{l\in L_{k}}\|\eta_{k}{\mathfrak{A}}^{\alpha}_{T,j}(\eta_{l}f)\|_{p_{n}}^{p_{n}}$
$\leq$ $C_{c_{0},\alpha,\delta}2^{-\varepsilon p_{n}j}\mathrm{e}^{-(n-1)T}\mathrm{e}^{(n-1)p_{n}T}\|f\|_{p_{n}}^{p_{n}},$
which proves (4.21). Here we use (4.23) in the second and the last inequalities, and the penultimate inequality follows from (4.22). Hence to prove (4.21) it suffices to show (4.22).
Let us now prove (4.22). To do this, we define another cutoff function $\widetilde{\phi}_{s}\in C_{c}^{\infty}(\mathbb{R})$ by $\widetilde{\phi}_{s}\equiv\phi_{s}$ on $[s,\infty)$ and $\widetilde{\phi}_{s}\equiv 0$ on $(-\infty,s-1]$. Note that in view of (4.12), we can write
(4.24) $\mathscr{A}^{\alpha}_{t,j}(r)=\widetilde{\phi}_{T}(r)\,\mathrm{e}^{-(n-1)(r+t)/2}\int_{\mathbb{R}}\beta_{j}(\lambda)\,\mathrm{e}^{\mathrm{i}\lambda(t-r)}{{\mathcal{H}}^{\alpha}(\lambda)\over{\mathcal{H}}(\lambda)}a_{2}^{\alpha}(\lambda,t)\,a_{2}^{0}(-\lambda,r)\,\mathrm{d}\lambda+\mathfrak{R}^{\alpha}_{t,j}(r),$
where
$\mathfrak{R}^{\alpha}_{t,j}(r)$ $=$ $\mathrm{e}^{-(n-1)(r+t)/2}\left((\phi_{T}(r)-\widetilde{\phi}_{T}(r)-\phi_{0}(r))\int_{\mathbb{R}}\beta_{j}(\lambda)\,\mathrm{e}^{\mathrm{i}\lambda(t-r)}{{\mathcal{H}}^{\alpha}(\lambda)\over{\mathcal{H}}(\lambda)}a_{2}^{\alpha}(\lambda,t)\,a_{2}^{0}(-\lambda,r)\,\mathrm{d}\lambda\right.$
$+\left.(\phi_{T}(r)-\phi_{0}(r))\int_{\mathbb{R}}\beta_{j}(\lambda)\,\mathrm{e}^{\mathrm{i}\lambda(t+r)}{{\mathcal{H}}^{\alpha}(\lambda)\over{\mathcal{H}}(-\lambda)}a_{2}^{\alpha}(\lambda,t)\,a_{2}^{0}(\lambda,r)\,\mathrm{d}\lambda\right).$
Lemmas 2.1, 2.2 and 2.4(iii) and integration by parts give that
$\sup_{t\in[T,T+1]}|\mathfrak{R}^{\alpha}_{t,j}(r)|\leq C_{\alpha,N}\mathrm{e}^{-(n-1)(r+t)/2}2^{-Nj}\phi_{T}(r)\,(1+|t-r|)^{-N}.$
Hence by Hölder’s inequality we have
$\left\|\sup_{t\in[T,T+1]}|(\psi_{0}(\psi_{1}f)*\mathfrak{R}^{\alpha}_{t,j})|\right\|_{p_{n}}\leq C_{c_{0},\alpha,N}2^{-Nj}\mathrm{e}^{-(n-1)t/2}\|\mathbbm{1}_{\mathop{\mathrm{supp}}\psi_{1}}f\|_{p_{n}},$
which indicates that, in view of (4.22), we are able to replace the convolution kernel $\mathscr{A}^{\alpha}_{t,j}(r)$ by another kernel
(4.25) $\widetilde{\mathscr{A}}^{\alpha}_{t,j}(z,w)=\psi_{0}(z)\,\psi_{1}(w)\,\mathrm{e}^{-(n-1)(d(z,w)+t)/2}\int_{\mathbb{R}}\beta_{j}(\lambda)\,\mathrm{e}^{\mathrm{i}\lambda(t-d(z,w))}b^{\alpha}(\lambda,t,d(z,w))\,\mathrm{d}\lambda$
with
(4.26) $b^{\alpha}(\lambda,t,r)=\widetilde{\phi}_{T}(r){{\mathcal{H}}^{\alpha}(\lambda)\over{\mathcal{H}}(\lambda)}a_{2}^{\alpha}(\lambda,t)\,a_{2}^{0}(-\lambda,r)\in S^{-\mathop{\mathrm{Re}}\alpha}_{\lambda}$
by Lemmas 2.2 and 2.4(iii). To show (4.22), it remains to show that
(4.27) $\left\|\sup_{t\in[T,T+1]}\left|\int_{\mathbb{H}^{n}}f(w)\,\widetilde{\mathscr{A}}^{\alpha}_{t,j}(\cdot,w)\,\mathrm{d}w\right|\right\|_{p_{n}}\leq C_{\alpha,\delta}2^{-\varepsilon j}\mathrm{e}^{-(n-1)T/p_{n}}\|\mathbbm{1}_{\mathop{\mathrm{supp}}\psi_{1}}f\|_{p_{n}}.$
The estimate (4.27) follows from the local smoothing estimate (4.7) in Proposition 4.3. To see this, we shall adopt the Iwasawa coordinates on $\mathbb{H}^{n}$ (see [10, p. 294]). Recall that $\mathbb{H}^{n}\cong\mathbb{R}^{n-1}\rtimes\mathbb{R}$ with multiplication law given by
$(v,u)\cdot(v^{\prime},u^{\prime})=(v+\mathrm{e}^{u}v^{\prime},u+u^{\prime}).$
Also under these coordinates the geodesic distance between $(v,u)$ and $(v^{\prime},u^{\prime})$ is given by
(4.28) $d((v,u),(v^{\prime},u^{\prime}))=\mathop{\mathrm{arcosh}}(\mathrm{e}^{-u-u^{\prime}}|v-v^{\prime}|^{2}+\cosh(u-u^{\prime})).$
Note that $d$ is invariant under left translations.
Without losing any generality, we assume the cutoff functions $\psi_{0},\psi_{1}$ in (4.25) satisfy $\mathop{\mathrm{supp}}\psi_{0}\subset B_{c_{0}}(0,0)$ and $\mathop{\mathrm{supp}}\psi_{1}\subset B_{c_{0}}(0,-T_{0})$, where $T_{0}\in[T-1-2c_{0},T+2+2c_{0}]$. Let $(v,u)\in\mathop{\mathrm{supp}}\psi_{0}$ and $(0,-T_{0})\cdot(e^{T_{0}}v^{\prime},u^{\prime})=(v^{\prime},u^{\prime}-T_{0})\in\mathop{\mathrm{supp}}\psi_{1}$. In view of (4.28) this implies $|v|,|u|,|u^{\prime}|\leq Cc_{0}$ and $|v^{\prime}|\leq Cc_{0}\,\mathrm{e}^{-T_{0}}$ with $C$ independent of $T_{0}$. Hence by (4.28) and the Taylor formula we can write
(4.29) $d((v,u),(v^{\prime},u^{\prime}-T_{0}))=T_{0}+u-u^{\prime}+(v-v^{\prime})^{2}A(v,u,v^{\prime},u^{\prime};T_{0}),$
where $A(\cdot,\cdot,\cdot,\cdot;T_{0})$ is a smooth function satisfying $A(0,0,0,0;T_{0})\sim 1$ uniformly in $T_{0}>5$, and that all the derivatives of $A(\cdot,\cdot,\cdot,\cdot;T_{0})$ at $(0,0,0,0)$ are uniformly bounded in $T_{0}>5$.
In view of (4.25), we set the function $\Phi((v,u),(v^{\prime},u^{\prime}))$ in (4.4) to be $d((v,u),(v^{\prime},u^{\prime}-T_{0}))$ and the symbol $a(\lambda,t,(v,u),(v^{\prime},u^{\prime}))$ in (4.4) to be $b^{\alpha}(\lambda,t,d((v,u),(v^{\prime},u^{\prime}-T_{0})))$. Notice that the problem in $[(v,u),(v^{\prime},u^{\prime})]$-coordinates is no longer degenerate. Indeed, on the one hand, by (4.29) the Monge-Ampere matrix of $\Phi$ turns into
$\left(\begin{array}[]{ccccc}0&\varepsilon_{1,2}&\dots&\varepsilon_{1,n-1}&1+\varepsilon_{1,n}\\ \varepsilon_{2,1}&-2a+\varepsilon_{2,2}&\dots&\varepsilon_{2,n-1}&\varepsilon_{2,n}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \varepsilon_{n-1,1}&\varepsilon_{n-1,2}&\dots&-2a+\varepsilon_{n-1,n-1}&\varepsilon_{n-1,n}\\ -1+\varepsilon_{n,1}&\varepsilon_{n,2}&\dots&\varepsilon_{n,n-1}&\varepsilon_{n,n}\end{array}\right),$
where $a=A(0,0,0,0;T_{0})$ and each $|\varepsilon_{i,j}|\leq Cc_{0}$ uniformly in $T_{0}>5$. Hence (4.1) is fulfilled provided $c_{0}$ is sufficiently small but independent of $T_{0}$. On the other hand, if we let $q$ be the homogeneous function in (4.2) satisfying $q^{2}((v,u),\xi)=\mathrm{e}^{2u}(\xi_{1}^{2}+\dots+\xi_{n-1}^{2})+\xi_{n}^{2}$, which is the cometric on $\mathbb{H}^{n}$, then for $u,v,u^{\prime},v^{\prime}$ as above one has
$q\!\left((v,u),{\partial d((v,u),(v^{\prime},u^{\prime}-T_{0}))\over\partial(v,u)}\right)=|(\nabla d(\cdot,(v^{\prime},u^{\prime}-T_{0})))(v,u)|_{g}\equiv 1,$
which is (4.2). Here $\nabla$ is the gradient and $|\cdot|_{g}$ is the norm both induced by the Riemannian metric on $\mathbb{H}^{n}$. Also, in view of (4.26) and (4.29), $b^{\alpha}(\lambda,t,d((v,u),(v^{\prime},u^{\prime}-T_{0})))$ satisfies (4.3) with $m$ replaced by $-\mathop{\mathrm{Re}}\alpha$. Hence by (4.7) in Proposition 4.3, we can write
$\int_{\mathbb{R}^{n}}\sup_{t\in[T,T+1]}\left|\int_{\mathbb{R}^{n}}f(v^{\prime},u^{\prime}-T_{0})\,\widetilde{\mathscr{A}}^{\alpha}_{t,j}((v,u),(v^{\prime},u^{\prime}-T_{0}))\,\mathrm{d}v^{\prime}\mathrm{d}u^{\prime}\right|^{p_{n}}\mathrm{d}v\mathrm{d}u$
$\qquad\leq C_{\alpha,\delta}2^{-\varepsilon p_{n}j}\mathrm{e}^{-(n-1)p_{n}T}\int_{\mathop{\mathrm{supp}}\psi_{1}}|f(v^{\prime},u^{\prime}-T_{0})|^{p_{n}}\,\mathrm{d}v^{\prime}\mathrm{d}u^{\prime},$
where the term $\mathrm{e}^{-(n-1)p_{n}T}$ comes from the exponential decay in (4.25). Note that under the Iwasawa coordinates, the measure induced by the Riemannian metric reads $\mathrm{e}^{-(n-1)u}\,\mathrm{d}v\mathrm{d}u$. So
$\left\|\sup_{t\in[T,T+1]}\left|\int_{\mathbb{H}^{n}}f(w)\,\widetilde{\mathscr{A}}^{\alpha}_{t,j}(\cdot,w)\,\mathrm{d}w\right|\right\|_{p_{n}}\leq C_{\alpha,\delta}2^{-\varepsilon j}\mathrm{e}^{-(n-1)T/p_{n}}\|\mathbbm{1}_{\mathop{\mathrm{supp}}\psi_{1}}f\|_{p_{n}},$
which is (4.27). ∎
Now we apply Lemma 4.7, Lemma 4.8 and Lemma 4.9 to prove Proposition 4.5.
Proof of Proposition 4.5.
Let $\delta$ be as in Lemma 4.9. By Lemmas 4.7 and 4.8 and complex interpolation, we have
(4.31) $\|\mathfrak{m}^{\alpha}_{T,j}(f)\|_{p_{n}}\leq C_{\alpha}2^{-(\mathop{\mathrm{Re}}\alpha+(n-2)/p_{n}-\delta)j}\mathrm{e}^{-(n-1)T/p_{n}}(1+T)^{2/p_{n}}\|f\|_{p_{n}}.$
In order to sum over $j$, we choose $j_{0}$ with $2^{j_{0}}=\mathrm{e}^{(n-1)p_{n}}(1+T)^{-2}$ such that the two norms in (4.20) and (4.31) are equal. The result is
$\|\mathfrak{m}^{\alpha}_{T}(f)\|_{p_{n}}$$\leq$ $\sum_{j=0}^{j_{0}}\|\mathfrak{m}^{\alpha}_{T,j}(f)\|_{p_{n}}+\sum_{j=j_{0}}^{\infty}\|\mathfrak{m}^{\alpha}_{T,j}(f)\|_{p_{n}}$
$\leq$ $C_{\alpha,\delta}\mathrm{e}^{[-(\mathop{\mathrm{Re}}\alpha+(n-1)/p_{n}-\delta)(n-1)p_{n}+(n-1)(1-1/p_{n})]T}(1+T)^{-2(\mathop{\mathrm{Re}}\alpha+(n-1)/p_{n}-\delta)}\|f\|_{p_{n}},$
which proves Proposition 4.5. ∎
Next we turn to handle the “local” part of the operator $\mathfrak{m}^{\alpha}$ by proving the following result.
Proposition 4.10.
Let $p_{n}$ be as in (1.14). If $\mathop{\mathrm{Re}}\alpha>(1-n)/p_{n}$, then
(4.32) $\left\|\sup_{0<t\leq 10}|M^{\alpha}_{t}(f)|\right\|_{p_{n}}\leq C_{\alpha}\|f\|_{p_{n}}.$
Proof.
Consider a partition of unity $\{\varkappa_{k}\}_{k\in\widetilde{K}}$ on $\mathbb{H}^{n}$ with each $\mathop{\mathrm{supp}}\varkappa_{k}$ contained in a ball $B_{k}$ of radius $10$, where the balls $B_{k}$ satisfy that $(1/5)B_{k}$ are mutually disjoint. Let $\varsigma_{k}\in C_{c}^{\infty}(\mathbb{H}^{n})$ such that $\varsigma_{k}\equiv 1$ on $2B_{k}$ and $\varsigma_{k}\equiv 0$ outside $3B_{k}$. By (1.7) we have that $M^{\alpha}_{t}(\varkappa_{k}f)=\varsigma_{k}M^{\alpha}_{t}(\varkappa_{k}f)$ for $0<t\leq 10$.
To prove Proposition 4.10, it is enough to show that for all $\alpha\in\mathbb{C}$ with $\mathop{\mathrm{Re}}\alpha>(1-n)/p_{n}$ there holds
(4.33) $\left\|\sup_{t\in[1,10]}|\psi_{0}\,M^{\alpha}_{t}(\psi_{1}f)|\right\|_{p_{n}}+\left\|\sup_{0<t\leq 1}|\psi_{0}\,M^{\alpha}_{t}(\psi_{1}f)|\right\|_{p_{n}}\leq C_{\alpha}\|\mathbbm{1}_{\mathop{\mathrm{supp}}\psi_{1}}f\|_{p_{n}}$
where $\psi_{0},\psi_{1}\in C^{\infty}_{c}(\mathbb{H}^{n})$ have sufficiently small supports.
Let us prove (4.33) for the case $1\leq t\leq 10$. Since the convolution kernel $K^{\alpha}_{t}$ of $M^{\alpha}_{t}$ satisfies $\mathop{\mathrm{supp}}K^{\alpha}_{t}(r)\subset\{r:0\leq r\leq t\}$ (see (1.6)), we write
$K^{\alpha}_{t}(r)=\phi_{t}(r)\int_{0}^{\infty}m^{\alpha}_{t}(\lambda)\,\varphi_{\lambda}(r)\,|{\mathcal{H}}(\lambda)|^{-2}\,\mathrm{d}\lambda=:\mathscr{B}^{\alpha}_{t}(r)+\mathscr{D}^{\alpha}_{t}(r),$
where by Lemma 2.1 and the asymptotics of $m^{\alpha}_{t}$ and $\varphi_{\lambda}$ (see Lemma 2.4(iii) and Remark 2.5)
(4.34) $\mathscr{B}^{\alpha}_{t}(r)=(\phi_{t}(r)-\phi_{0}(r))\,\mathrm{e}^{-(n-1)(r+t)/2}\int_{\mathbb{R}}\mathrm{e}^{\mathrm{i}\lambda(t-r)}{{\mathcal{H}}^{\alpha}(\lambda)\over{\mathcal{H}}(\lambda)}a^{\alpha}_{2}(\lambda,t)\,a^{0}_{2}(-\lambda,r)\,\mathrm{d}\lambda,$
$\mathscr{D}^{\alpha}_{t}(r)=K^{\alpha}_{t}(r)-\mathscr{B}^{\alpha}_{t}(r)$ and $\phi_{s}$ is the cutoff function as above.
Observe that in terms of (4.33), $\mathscr{D}^{\alpha}_{t}$ is an error term. To see this, by Lemmas 2.1, 2.2 and 2.4 together with an argument in the proof of Lemma 4.6, we have
(4.35) $|\mathscr{D}^{\alpha}_{t}(r)|\leq C_{\alpha}(\phi_{10}(r)+\phi_{0}(r)\,r^{\mathop{\mathrm{Re}}\alpha-(n+1)/2}),$
which implies
$\sup_{z\in\mathbb{H}^{n}}\left\|\sup_{t\in[1,10]}|\mathscr{D}^{\alpha}_{t}(d(z,\cdot))|\right\|_{1}+\sup_{w\in\mathbb{H}^{n}}\left\|\sup_{t\in[1,10]}|\mathscr{D}^{\alpha}_{t}(d(\cdot,w))|\right\|_{1}\leq C_{\alpha}.$
By Schur’s lemma, we obtain
(4.36) $\left\|\sup_{t\in[1,10]}|\psi_{0}\,((\psi_{1}f)*\mathscr{D}^{\alpha}_{t})|\right\|_{p_{n}}\leq C_{\alpha}\|f\|_{p_{n}}.$
Let us turn to the term $\mathscr{B}^{\alpha}_{t}$. In view of the remark following [20, Corollary 2.2], under local coordinates, for $1\leq t\leq 10$, if we set the function $\Phi$ in (4.4) to be $d(z,w)$ and the symbol $a$ in (4.4) to be
$(\phi_{t}(d(z,w))-\phi_{0}(d(z,w))){{\mathcal{H}}^{\alpha}(\lambda)\over{\mathcal{H}}(\lambda)}a^{\alpha}_{2}(\lambda,t)\,a^{0}_{2}(-\lambda,d(z,w)),$
then the conditions on $\Phi$ and $a$ in Proposition 4.3 are fulfilled. Therefore we can invoke (4.8) in Proposition 4.3 to obtain
$\left\|\sup_{t\in[1,10]}|\psi_{0}\,((\psi_{1}f)*\mathscr{B}^{\alpha}_{t})|\right\|_{p_{n}}\leq C_{\alpha}\|\mathbbm{1}_{\mathop{\mathrm{supp}}\psi_{1}}f\|_{p_{n}}\ \ \ \ \mbox{for }\mathop{\mathrm{Re}}\alpha>{1-n\over p_{n}}.$
This, combined with (4.36), concludes the proof of (4.33) for the case $1\leq t\leq 10$.
Now we consider (4.33) for the case $0<t\leq 1$. Let $\gamma$ be the cutoff function in the proof of Lemma 4.6. We define another cutoff function $\widetilde{\gamma}\in C_{c}^{\infty}(\mathbb{R}_{+})$ with $\widetilde{\gamma}\equiv 1$ on $[1/2,\infty)$ and $\widetilde{\gamma}\equiv 0$ on $[0,1/4]$. Let $K^{\alpha}_{t}(r)$ be the convolution kernel of $M^{\alpha}_{t}$. Note that in view of (1.7), $\mathop{\mathrm{supp}}K^{\alpha}_{t}(r)\subset\{r:0\leq r\leq t\}$. So we can write
$K^{\alpha}_{t}(r)=\gamma\!\left({r\over t}\right)\int_{0}^{\infty}m^{\alpha}_{t}(\lambda)\,\varphi_{\lambda}(r)\,|{\mathcal{H}}(\lambda)|^{-2}\,\mathrm{d}\lambda=\mathscr{P}^{\alpha}_{t}(r)+\mathscr{W}^{\alpha}_{t}(r),$
where by the asymptotics of $m^{\alpha}_{t}$ and $\varphi_{\lambda}$ (see Lemma 2.4(ii) and Remark 2.5)
$\mathscr{P}^{\alpha}_{t}(r)=\widetilde{\gamma}\!\left({r\over t}\right)\int_{\mathbb{R}}\mathrm{e}^{\mathrm{i}\lambda(t-r)}a^{\alpha,N}_{1}(\lambda,t)\,a^{0,N}_{1}(-\lambda,r)\,(1-\gamma(\lambda r))\,|{\mathcal{H}}(\lambda)|^{-2}\,\mathrm{d}\lambda$
and $\mathscr{W}^{\alpha}_{t}(r)=K^{\alpha}_{t}(r)-\mathscr{P}^{\alpha}_{t}(r)$.
Observe that in terms of (4.33), $\mathscr{W}^{\alpha}_{t}$ is an error term. To see this, by Lemmas 2.1, 2.2 and 2.4, an argument in the proof of Lemma 4.6 shows that
$|\mathscr{W}^{\alpha}_{t}(r)|\leq C_{\alpha}t^{-n}\mathbbm{1}_{[0,2t]}(r)\left(1+\left({t\over r}\right)^{-\mathop{\mathrm{Re}}\alpha+(n+1)/2}\right),$
and so
(4.37) $\left\|\sup_{0<t\leq 1}|\psi_{0}\,((\psi_{1}f)*\mathscr{W}^{\alpha}_{t})|\right\|_{p_{n}}\leq C_{\alpha}\|f\|_{p_{n}}\ \ \ \mbox{for }\mathop{\mathrm{Re}}\alpha>{1-n\over p_{n}}.$
Now we claim that for $\mathop{\mathrm{Re}}\alpha>{(1-n)/p_{n}}$,
(4.38) $\left\|\sup_{0<t\leq 1}|\psi_{0}\,((\psi_{1}f)*\mathscr{P}^{\alpha}_{t})|\right\|_{p_{n}}\leq C_{\alpha}\|f\|_{p_{n}}.$
To prove (4.38), we observe that
$\mathscr{P}^{\alpha}_{t}(r)=\widetilde{\widetilde{\gamma}}\!\left({r\over t}\right)t^{-n}\int_{\mathbb{R}}\mathrm{e}^{\mathrm{i}\lambda(1-r/t)}|\lambda|^{-\mathop{\mathrm{Re}}\alpha}\widetilde{a}^{\alpha}_{1}(\lambda,t)\,\widetilde{a}^{0}_{1}\!\left(\lambda{r\over t},r\right)\left(1-\gamma\!\left(\lambda{r\over t}\right)\right)\mathfrak{w}\!\left({\lambda\over t}\right)\mathrm{d}\lambda,$
where $\widetilde{\widetilde{\gamma}}(s)=s\,\widetilde{\gamma}(s)$, $\mathfrak{w}(\lambda)=|\lambda|^{-(n-1)}|{\mathcal{H}}(\lambda)|^{-2}$ and $\widetilde{a}^{\alpha}_{1}(\lambda,t)=|\lambda|^{-\mathop{\mathrm{Re}}\alpha-(n-1)/2}a^{\alpha,N}_{1}(\lambda/t,t)$. By Lemma 2.2 and (2.13), we see that $\mathfrak{w}(\lambda),\widetilde{a}^{\alpha}_{1}(\lambda,t)\in S^{0}_{\lambda}$. Without losing any generality, we assume the supports of $\psi_{0}$ and $\psi_{1}$ in (4.38) are in some neighborhoods of $\mathbf{0}$. Here we adopt the hyperboloid coordinates. Let
$\begin{array}[]{cccc}\tau:&\mathbb{R}^{n}&\to&\mathbb{H}^{n},\\ &x&\mapsto&(\sqrt{1+|x|^{2}},x).\end{array}$
Under this coordinates, the geodesic distance reads
$d(\tau(x),\tau(y))=\mathop{\mathrm{arcosh}}(\sqrt{1+|x|^{2}}\sqrt{1+|y|^{2}}-x\cdot y).$
We set $\widetilde{d}^{t}(x,y)=t^{-1}d(t\,\tau(x),t\,\tau(y))$ and
$\widetilde{\mathscr{P}}^{\alpha}_{t}(r)=t^{n}\mathscr{P}^{\alpha}_{t}(tr)=\widetilde{\widetilde{\gamma}}(r)\int_{\mathbb{R}}\mathrm{e}^{\mathrm{i}\lambda(1-r)}|\lambda|^{-\mathop{\mathrm{Re}}\alpha}\widetilde{a}^{\alpha}_{1}(\lambda,t)\,\widetilde{a}^{0}_{1}(\lambda r,tr)\,(1-\gamma(\lambda r))\,\mathfrak{w}\!\left({\lambda\over t}\right)\mathrm{d}\lambda.$
Then (4.38) is equivalent to
(4.40) $\left\|\sup_{0<t\leq 1}\left|\widetilde{\psi}_{0}(\cdot)\int_{\mathbb{R}^{n}}\widetilde{\psi}_{1}(y)\,g(y)\,\widetilde{\mathscr{P}}^{\alpha}_{t}(\widetilde{d}^{t}(\cdot,y))\,\mathrm{d}y\right|\right\|_{L^{p_{n}}(\mathbb{R}^{n})}\leq C_{\alpha}\|g\|_{L^{p_{n}}(\mathbb{R}^{n})}$
for $\mathop{\mathrm{Re}}\alpha>{(1-n)/p_{n}}.$
To prove (4.40), we set
$b(\lambda,t,x,y)=\widetilde{\widetilde{\gamma}}(\widetilde{d}^{t}(x,y))\,(1-\gamma(\lambda\widetilde{d}^{t}(x,y)))\,|\lambda|^{-\mathop{\mathrm{Re}}\alpha}\widetilde{a}^{\alpha}_{1}(\lambda,t)\,\widetilde{a}^{0}_{1}(\lambda\widetilde{d}^{t}(x,y),t\widetilde{d}^{t}(x,y))\,\mathfrak{w}\!\left({\lambda\over t}\right).$
In view of [21, Proposition 6.2.4], modulo a $C^{\infty}$ error, for $0<t\leq 1$, $\mathscr{P}^{\alpha}_{t}(\widetilde{d}^{t}(x,y))$ can be written as the sum of finite terms, each of which is of the form
(4.41) $\widetilde{\widetilde{\gamma}}(\widetilde{d}^{t}(x,y))\,\widetilde{\widetilde{\psi}}_{0}(x)\,\widetilde{\widetilde{\psi}}_{1}(y)\int_{\mathbb{R}^{n}}\mathrm{e}^{\mathrm{i}(\widetilde{\varphi}(x,t,\xi)-y\cdot\xi)}\widetilde{b}(x,t,\xi)\,\mathrm{d}\xi,$
where
$|\partial^{\gamma_{1}}_{x}\partial^{\gamma_{2}}_{\xi}\widetilde{b}(x,t,\xi)|\leq C_{\alpha,\gamma_{1},\gamma_{2}}(1+|\xi|)^{-\mathop{\mathrm{Re}}\alpha-(n-1)/2-|\gamma_{2}|}\qquad\mbox{for }0<t\leq 2^{-j_{0}},$
$|\partial^{\gamma_{1}}_{x}\partial^{k}_{t}\partial^{\gamma_{2}}_{\xi}\widetilde{b}(x,t,\xi)|\leq C_{\alpha,k,\gamma_{1},\gamma_{2}}(1+|\xi|)^{-\mathop{\mathrm{Re}}\alpha-(n-1)/2-|\gamma_{2}|}\qquad\mbox{for }t\in[2^{-j_{0}},1]$
for an integer $j_{0}$ to be chosen later, and $\widetilde{\varphi}$ is homogeneous of degree $1$ on $\xi$ satisfying $\det({\partial^{2}\widetilde{\varphi}/\partial x\partial\xi})\neq 0$. If we set $T_{t}\ (0<t\leq 1)$ to be the operator given by integration against the Schwartz kernel (4.41), then by [1, Proposition 3.2] and [5, Theorem 1.4],
$\|T_{t}(g)(x)\|_{L^{p_{n}}_{x,t}(\mathbb{R}^{n}\times[2^{-j_{0}},1])}\leq C_{\alpha}\|g\|_{W^{-\mathop{\mathrm{Re}}\alpha-(n-1)/p_{n}-s,p_{n}}(\mathbb{R}^{n})}\qquad\mbox{for any }s<{1\over p_{n}}.$
We then use an argument in the proof of [21, Theorem 6.3.1] to obtain $\|\sup_{t\in[2^{-j_{0}},1]}|T_{t}(g)|\|_{p_{n}}\leq C_{\alpha}\|g\|_{p_{n}}.$ Hence, the proof of (4.40) reduces to showing
(4.42) $\left\|\sup_{0<t\leq 2^{-j_{0}}}|T_{t}(g)|\right\|_{p_{n}}\leq C_{\alpha}\|g\|_{p_{n}}.$
For this, we need to consider $\mathscr{P}^{\alpha}_{t}(\widetilde{d}^{t}(x,y))$ when $t\to 0$. Note that $\widetilde{d}^{t}(x,y)\to|x-y|$ as $t\to 0$ uniformly in $x,y$ belonging to a fixed compact set. We also have that $\mathfrak{w}(\lambda)\to|C_{n}^{\prime}|^{-2}$ as $\lambda\to\infty$, with $C_{n}^{\prime}=2^{n-2}\Gamma(n/2)/\sqrt{\uppi}$ being the constant in (2.3). It tells us that $\widetilde{K}^{\alpha}_{0}(\widetilde{d}^{0}(x,y))$ is of the form
$\widetilde{\widetilde{\gamma}}(|x-y|)\int_{\mathbb{R}}\mathrm{e}^{\mathrm{i}\lambda(1-|x-y|)}|\lambda|^{-\mathop{\mathrm{Re}}\alpha}\widetilde{a}^{\alpha}_{1}(\lambda,0)\,\widetilde{a}^{0}_{1}(\lambda|x-y|,0)\,(1-\gamma(\lambda|x-y|))\,\mathrm{d}\lambda.$
This, together with [21, Proposition 6.2.4] and the asymptotics of the Fourier transform of the spherical measure on $\mathbb{R}^{n}$ derived using [25, Chapter VIII, (15) and (25)], indicates that $\widetilde{\varphi}$ in (4.41) satisfies $\widetilde{\varphi}(x,0,\xi)=x\cdot\xi\pm|\xi|$. Choosing $j_{0}$ sufficiently large and noting that $\widetilde{\varphi}(x,t,\cdot)$ is homogeneous of degree $1$, we use Taylor’s formula to derive that the phase function in (4.41) satisfies
$\widetilde{\varphi}(x,t,\xi)=x\cdot\xi\pm|\xi|+O(t|\xi|)\qquad\mbox{for }0<t\leq 2^{-j_{0}}.$
Then we rescale the integral representation of $T_{t}$ into the form of the operator $\mathcal{F}_{t}$ on [21, p. 192]. Using the argument showing [21, Corollary 6.3.3], together with [1, Proposition 3.2] and [5, Theorem 1.4], we obtain (4.42), and the estimate (4.40) is valid. This proves (4.38) and finishes the proof of (4.33) for the case $0<t\leq 1$. Hence, the proof of Proposition 4.10 is concluded. ∎
Proof of Theorem 4.4.
This is a direct consequence of (4.10) and Proposition 4.10. ∎
We finally come to the
Proof of Theorem 1.2.
We shall use the interpolation argument in [13]. In view of [13, Theorem 3], $\mathfrak{m}^{\alpha}$ is bounded on $L^{p}(\mathbb{H}^{n})$ for $\mathop{\mathrm{Re}}\alpha>0$ and $p=\infty$, or $\mathop{\mathrm{Re}}\alpha>(2-n)/2$ and $p=2$. Then we interpolate between these results and Theorem 4.4 to derive Theorem 1.2. ∎
Acknowledgements: The authors were supported by National Key R&\&&D Program of China 2022YFA1005700. P. Chen was supported by NNSF of China 12171489. The authors thank X. Chen, N.J. Liu, L. Roncal, L. Song and H.W. Zhang for helpful discussions and suggestions.